### v1

# Bilateral facial reduction: qualification-free subdifferential calculus and exact duality in convex analysis

Abstract
This paper introduces a geometric framework to extend convex analysis results beyond standard qualification conditions such as intersecting relative interiors of domains. We define the joint facial subspace $T$ as the span of the face of CDC-D generated by 0, with its affine translation $T_{a}$ containing $C\cap D$. The intersections $C\cap T_{a}$ and $D\cap T_{a}$ are faces of the original sets, establishing $T_{a}$ as a bilateral facial reduction with parallels to Borwein-Wolkowicz facial reduction in conic programming. When reducing the domains of two convex functions to their joint facial subspace $T_{a}$, their relative interiors then always intersect, enabling unqualified application of classical theorems via localization. Key generalizations include subdifferential additivity, normal cones of intersections, a subdifferential chain rule, attained infimal convolution for $(f+g)^{*},$ and an exact Fenchel-Rockafellar dual. We characterize $T$ via generated faces and computationally through an iterative process converging in at most $n$ steps in $\mathbb{R}^{n}$. Proofs are self-contained and introduce novel concepts like facial subspaces and nested normals to describe convex boundaries and the lattice of faces.
1 Introduction
In convex analysis, the effective domain of the sum of two convex functions is the intersection of their effective domains. Many central results, including the summation rule of subdifferentials, the convex conjugate of sums of functions, or the strong duality of the Fenchel-Rockafellar dual, exclude the case where the domains of the convex functions intersect only on their relative boundaries. The conditions excluding such cases are called qualification conditions, or constraint qualifications depending on the context. One such condition, widespread in the seminal work [14] by Rockafellar, is that for two functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$,
$\operatorname{relint}(\operatorname{dom}(f))\cap\operatorname{relint}(\operatorname{dom}(g))\neq\emptyset.$
For instance, this condition appears in [14, Theorem 23.8], which establishes the summation rule of subdifferentials:
$\partial(f+g)(x)=\partial f(x)+\partial g(x).$
A second example is the strong duality of the Fenchel-Rockafellar dual: for $A\in\mathbb{R}^{m\times n}$, denoting by $f^{*}$ the convex conjugate of $f$,
$\inf_{x\in\mathbb{R}^{n}}f(x)+g(Ax)=-\inf_{y\in\mathbb{R}^{m}}[f^{*}(A^{T}y)+g^{*}(-y)].$
In [14, Corollary 31.2.1], this is guaranteed to hold if
$\operatorname{relint}(\operatorname{dom}(f))\cap A^{-1}(\operatorname{relint}\operatorname{dom}(g))\neq\emptyset,$
where $A^{-1}(\cdot)$ is the preimage for the linear map induced by the matrix $A$. This requirement is often called a constraint qualification, and is strikingly similar to the qualification condition for the subdifferential sum rule.
These and a number of other results share similar conditions. The object of this work is to generalize such results beyond the regime where these conditions hold. Specifically, in Section 5 we find generalizations for the two aforementioned results, as well as a few others: the attainment of infimal convolution as the convex conjugate of the sum of two convex functions, the normal cone of the intersection of two convex sets, and a chain rule for the subdifferential. There are many reasons such generalizations are worthwhile, and we list two here. From a theoretical perspective, extending such results makes the theory more elegant and intuitive. A second motivation is that such generalizations could lead to improved methods in optimization, as did facial reduction in conical programming (see [3]).
Qualification conditions are not merely theoretical artifacts; they are supported by counter examples, which indicates that these conditions reflect more or less accurately the true scope of the statements. We consider one such counter-example for the sum rule of subdifferentials. Take the function $f(x)=-\sqrt{x}$ for $x\geq 0$ and $f(x)=\infty$ otherwise. This function has effective domain $\mathbb{R}_{+}$. Adding it to the indicator function $g(x)=\iota_{\mathbb{R}_{-}}(x)$, the sum rule states that $\partial(f+g)(0)=\partial f(0)+\partial g(0)=\emptyset+\mathbb{R}_{+}=\emptyset$, when in fact, $\partial(f+g)(0)=\partial\iota_{0}(0)=\mathbb{R}$. The relative interiors of the domains are $(-\infty,0)$ and $(0,\infty)$, and do not intersect. The qualification condition does not hold, so there is no contradiction with [14, Theorem 23.8], yet this still shows that the sum rule does not in fact apply to the sum of any two convex functions. Thus, an accurate generalization of the sum rule must offer a different formula for the subdifferential of the sum of two convex functions, and such a formula should ideally reduce to the classical sum rule when qualification conditions are satisfied. We provide such a formula in Theorem 5.1.
Our approach to relaxing qualification conditions is to focus on the geometry of the effective domain of the convex functions. Qualification conditions fail when relative interiors do not intersect, in which case only the boundaries of the domains meet. We show that the boundaries of two convex sets always meet in a structured way: each set presents a convex face to the other, and both of these faces are revealed by the intersection with a common affine subspace, which we name the joint facial subspace $T_{a}$. The main property of the joint facial subspace is that given convex sets $C,D\subseteq\mathbb{R}^{n}$, with $T_{a}$ defined as in Definition 4.1, whenever $C\cap D\neq\emptyset$,
$\operatorname{relint}(C\cap T_{a})\cap\operatorname{relint}(D\cap T_{a})\neq\emptyset.$
The results which shows this is Corollary 4.3. Therefore, qualification conditions always hold when localizing to the joint facial subspace. Roughly, we generalize results in convex analysis by replacing convex functions $f$ by their localized version $f+\iota_{T_{a}}$. The modified results then hold even without qualification conditions. And when qualification conditions do hold for the original results, then $T_{a}=\mathbb{R}^{n}$ (see Proposition 5.1), and the additional indicator function can be ignored.
For instance, in Theorem 5.1 we state that
$\partial(f+g)(x)=\partial(f+\iota_{T_{a}})(x)+\partial(g+\iota_{T_{a}})(x),$
without any condition on $f$ and $g$ other than they be convex functions. This modification works for the previously encountered counter example. In that problem, where $f(x)=-\sqrt{x}$ and $g(x)=\iota_{(-\infty,0]}$, the joint facial subspace from Definition 4.1 is $T_{a}=\{0\}$, and so
$\partial(f+\iota_{T_{a}})(x)+\partial(g+\iota_{T_{a}})(x)=\partial\iota_{\{0\}}+\partial\iota_{\{0\}}=\mathbb{R},$
which is indeed $\partial(f+g)(x)$. One way to interpret our result is that there are subgradients that are “hiding” in the faces of the domains of the convex functions. Intersecting with the joint facial subspace reveals these hidden gradients.
The central theoretical concept introduced in this work is the joint facial subspace. We therefore focus much of our efforts in describing this object. This is the focus of Section 4, where we define the joint facial subspace in Definition 4.1 , provide two characterizations and state core property, that qualification conditions hold when localizing to it. Given two convex sets $C,D\subseteq\mathbb{R}^{n}$, the joint facial subspace is defined from the Minkowski difference $C-D=\{c-d\mid c\in C,d\in D\}$. In the first characterization, Theorem 4.1, it is made clear how the joint facial subspace relates to $C$ and $D$ individually. The third characterization is computational in nature, and describes how the joint facial subspace can be found from a number of steps that is bounded by the ambient dimension.
Another important contribution of the present work is in our proof techniques. In Section 6, we build a theoretical framework with nested normals (Definition 6.3) and facial subspaces (Definition 6.2), two objects which together describe the lattice of faces of a convex set. Their combination forms a dual object which is more descriptive than the standard normal cone (see Theorem 6.2). Our theory is therefore at least in part a theory about the boundary of convex sets. From this toolkit, we are able to describe the faces of differences of convex sets CDC-D in Section 6.2, from which we derive the various properties of the joint facial subspace in Section 4. These objects will be further explored and refined in future work.
The structure of the paper goes as follows. In Section 2 surveys prior works and Section 3 introduces some notation. Section 4 is as described above, with proofs deferred to Section 7. In Section 5, we provide generalizations of central results in convex analysis, together with their proofs, which use Corollary 4.3 from Section 4. In Section 6, we give a motivated introduction to our mathematical framework, introducing key concepts. In Section 7, we prove the results in Section 4 by using tools from Section 6.
2 Prior works
The existing research which is the most related to this work is Facial Reduction (FR) in conic programming  [2, 1, 3], which helped guide the development of the theory in this work. We briefly describe this method. Given a feasible cone and affine constraints, Slater’s condition (that the interior of the feasible cone has at least one point satisfying the constraints), FR narrows the ambient space of the problem to the span of a “minimal” face of the feasible cone. After narrowing the problem via this “preprocessing” step, the problem is guaranteed to satisfy Slater’s condition.
Our joint facial subspace does something similar, but for any two intersecting convex sets. We find a “minimal” face of the difference of the sets, CDC-D (see Definition 4.1), and this yields the joint facial subspace. The computation of the joint facial subspace can be similarly viewed as a processing step, after which the problem is reduced to the joint facial subspace. As specified in [3], FR can be applied with finite-step algorithms, with a number of steps bounded by the ambient dimension. This is very similar to our second characterization of the joint facial subspace, Theorem 7.1.
A result which is related (where the relation is examined in [11]), is the Extended Dual System of Ramana, a dual program for semidefinite programming for which strong duality always holds [13]. The corresponding result in our work is the exact Fenchel-Rockafellar dual in Theorem 5.5.
From these strong parallels, we conjecture that many aspects of Facial Reduction and the Extended Dual System can be recovered—and potentially generalized—from results in this work.
Prior efforts to obtain a qualification-free sum rule for subdifferentials have developed powerful limiting approaches that achieve exact characterizations without traditional qualifications. For instance, $\epsilon$-subdifferentials define approximate subgradients satisfying the subgradient inequality up to an $\varepsilon$-error, recovering the exact subdifferential as $\varepsilon$ approaches zero [4, 12, 10]. Similarly, fuzzy sum rules approximate the subdifferential by perturbing the point $x$ at which the subdifferential is taken, deriving exact additivity rules in the limit of vanishing perturbations [6, 5]. These methods, leveraging local neighborhood behavior, are related to ideas in variational analysis, and offer versatile tools for theoretical and computational analysis across general spaces. They can often be implemented with fixed tolerances for practical computation, though the exactness relies on infinite limiting processes to eliminate approximation errors. Though our approach is tailored to pairs of functions, our formula for the subdifferential of a sum of convex functions is exact, and can be computed in finite steps.
One theoretical insight that was critical to developing our theory was the lexicographical characterization of facial subspaces from [8]. This work shows that every face of a convex set can be obtained by iteratively exposing faces. It is a sequel of a preceding result [9], which shows a separation theorem for convex sets. Though the lexicographical framework was not flexible enough for our needs, prompting us to develop an independent theory, we did draw inspiration from the results and the proofs. In this paper, we recover this lexicographical representation, though in different terms, primarily in Lemma 6.5. Our proof is direct, and does not make use of a separation theorem.
3 Preliminaries
For two sets $A,B\subseteq\mathbb{R}^{n}$, $A+B:=\{a+b\mid a\in A,b\in B\}$ is the Minkowski sum, with $A-B:=\{a-b\mid a\in A,b\in B\}$, and for $z\in\mathbb{R}^{n}$, let $A+z:=A+\{z\}$. For a set $S\subseteq\mathbb{R}^{n}$, we denote by $\iota_{S}(x)$ the indicator function which is 0 if $x\in S$ and $\infty$ otherwise.
For a set $S\subseteq\mathbb{R}^{n}$, denote by $\operatorname{conv}(S)$ its convex hull. For $y_{1},y_{2}\in\mathbb{R}^{n}$, let $y_{1},y_{2}]:=\operatorname{conv}\{y_{1},y_{2}\}$, and $y_{1},y_{2}[:=\operatorname{conv}\{y_{1},y_{2}\}\setminus\{y_{1},y_{2}\}$.
Intuitively the notion of a face of a convex set $C$ is a subset of $C$ for which there is no line segment in $C$ that “goes through” $F$. The following definitions can be found inline in [16, p. 16] and [16, p.75] respectively.
Definition 3.1 (Face of a convex set).
Let $C$ be convex. A subset $F\subseteq C$ is defined to be a face of $C$ when
$\forall y,z\in C,(\exists x\in]y,z[\cap F\implies y,z\in F).$
Definition 3.2 (Exposed face of a convex set).
Let $C\subseteq\mathbb{R}^{n}$ be convex. A subset $F\subseteq C$ is defined to be an exposed face of $C$ when
$\exists u\in\mathbb{R}^{n}\setminus\{0\}:F=\operatorname*{\arg\max}_{c\in C}\langle c,u\rangle.$
For any set $S\subseteq C$, we denote by $F_{C}(S)$ the face of $C$ generated by $S$; it is the smallest face of $C$ that contains $S$ (it is also the intersection of all faces containing $S$, which, it can be shown, is itself a face).
We also recall some objects from convex analysis. Let $\bar{\mathbb{R}}=\mathbb{R}\cap\{-\infty,\infty\}$. For a convex function $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the subdifferential is
$\partial f(x):=\{v\in\mathbb{R}^{n}\mid\forall y\in\mathbb{R}^{n},\langle v,y-x\rangle\leq f(y)\},$
and the convex conjugate is $f^{*}(y):=\sup_{x\in\mathbb{R}^{n}}\langle y,x\rangle-f(x)$.
4 The joint facial subspace
In this section we investigate the following object.
Definition 4.1 (The joint facial subspace).
The joint facial subspace of two convex sets $C,D\subseteq\mathbb{R}^{n}$ is
$T(C,D):=\operatorname{\mathrm{span}}F_{C-D}(0).$
We also define $T_{a}(C,D):=T(C,D)+C\cap D$, the joint affine facial subspace.
The Minkowski sum $(+C\cap D)$ in the definition of $T_{a}$ is in fact a simple translation. This follows from the next corollary.
Corollary 4.1 (Containment of the intersection).
Let $C,D\in\mathbb{R}^{n}$ be convex. Then
$C\cap D-C\cap D\subseteq T(C,D).$
This result follows from Lemma 6.10; we defer the proof of Corollary 4.1 to the end of Section 6.2.
From Corollary 4.1, we see that picking any $x\in C\cap D$, we have $T_{a}=T+x$, independently of the choice of $x$. Corollary 4.1 also implies that $C\cap D\subseteq T_{a}$. The affine subspace $T_{a}$ is therefore the translation of $T$ that contains $C\cap D$.
In this section we provide two characterizations of the joint facial subspace. The next result shows the relation of the joint facial subspace to the individual convex sets, and builds towards the first characterization.
Proposition 4.1 (The joint facial subspace reveals faces).
For convex sets $C,D\subseteq\mathbb{R}^{n}$, and $T_{a}=T_{a}(C,D)$,
$T_{a}\cap C=F_{C}(C\cap D).$
For the first characterization we require a subspace that is related to $F_{C}(C\cap D)$. This subspace, the generated facial subspace, is central to our theory, as is made clear in Section 6. There, we will re-introduce this object in a motivated way, in Definition 6.5.
Definition 4.2 (Preliminary definition of the generated facial subspace).
Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, define the facial subspace of $C$ generated by $S\subseteq C$ to be
$H_{C}(S):=\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S)).$
The next theorem is our first characterization of the joint facial subspace.
Theorem 4.1 (Characterization through generated faces).
The joint facial subspace of $C,D\subseteq\mathbb{R}^{n}$ is
$T(C,D)=H_{C}(C\cap D)+H_{D}(C\cap D).$
Definition 4.1 and Theorem 4.1 both specify the joint facial subspace via analytical formulas. Our second characterization instead specifies the joint facial subspace as the result of a process with a number of steps bounded by the ambient dimension, which is more amenable to computation. This result, Theorem 7.1, is best stated in terms of nested normals a concept that we introduce in Definition 6.3, and motivate throughout Section 6. For this reason, we defer it to Section 7.3. For the moment, we state a direct corollary which does not require this object.
Corollary 4.2 (Iterative bilateral facial reduction).
Let $C,D\subseteq\mathbb{R}^{n}$ be convex sets and fix any $x\in C\cap D$. Beginning with $T_{0}=\mathbb{R}^{n}$, let
$T_{i+1}=T_{i}\cap(N_{C}(x)\cap[-N_{D}(x)]\cap T_{i})^{\perp}.$
Let $\ell$ be the smallest number such that $T_{\ell+1}=T_{\ell}$. Then:
  1. 1.
    $T_{\ell}=T(C,D)$.
  2. 2.
    $\ell\leq n$.
Now that the joint facial subspace has been thoroughly identified, we move on to describing its core property. Recall that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, a common qualification condition is that for convex sets $C:=\operatorname{dom}(f),D:=\operatorname{dom}(g)$,
$\operatorname{relint}(C)\cap\operatorname{relint}(D)\neq\emptyset.$ (1)
It was shown in [14, theorem 11.3] that this notion is equivalent to a lack of proper separation of the domains. We define the notions of separation and proper separation below.
Definition 4.3 (Separation of sets).
Two convex sets $C,D\subseteq\mathbb{R}^{n}$ are said to be separated if there exists some $v\in\mathbb{R}^{n}\setminus\{0\}$ such that
$\forall c\in C,d\in D,v\cdot c\leq v\cdot d.$ (2)
The concept of proper separation is slightly stronger.
Definition 4.4 (Proper separation of sets).
Sets $C,D\in\mathbb{R}^{n}$ are said to be properly separated when there exists some $v\in\mathbb{R}^{n}$ such that Equation 2 holds, and such that there is some pair $(c,d)\in C\times D$ for which the inequality in Equation 2 is strict.
Proper separation excludes the case where the two sets are merely separated by virtue of being both contained in some affine hyperplane.
We will show that inside the joint facial subspace, we always have a “lack of separation” of the convex sets. Since this regularity condition is stronger than a “lack of proper separation”, it is stronger than Equation 1.
Theorem 4.2 (Qualification conditions always holds in the joint facial subspace).
For two convex sets $C,D\subseteq\mathbb{R}^{n}$, $C\cap T_{a}(C,D)$ and $D\cap T_{a}(C,D)$ are not separated.
The proof of Theorem 4.2 is the subject of Section 7.1. The next corollary follows immediately from Theorem 4.2.
Corollary 4.3 (rint qualification holds in joint facial subspace).
For convex sets $C,D\subseteq\mathbb{R}^{n}$,
$\operatorname{relint}(C\cap T_{a})\cap\operatorname{relint}(D\cap T_{a})\neq\emptyset.$
With Corollary 4.3 we can use standard results of convex analysis from [14] by first taking intersections with $T_{a}$. We do this in the proof of each of the results in Section 5. Also note that if we only had access to weaker results than those available in [14], results that would have the stronger assumption of “no separation of the sets” instead of “intersection of relative interiors is non-empty”, the theory in this paper would still hold, and indeed we would recover the stronger results with the weaker assumption, that the intersection of relative interiors is non-empty.
5 Main results
In this section we show five generalizations of important results of convex analysis: additivity of subdifferentials, normal cone of the intersection of two convex sets, attainment of the infimal convolution in the conjugate of a sum of two functions, the chain rule of the subdifferential, and the Fenchel-Rockafellar dual. All proofs are deferred to Section 7, because they require the machinery introduced in Section 6.
Proposition 5.1 (Recovery of the classical results).
In all results that follow, the classical analogue is recovered when qualification conditions hold. Indeed, given two convex functions $f,g$ such that their relative interiors intersect, we have $\operatorname{dom}(f),\operatorname{dom}(g)\subseteq T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$, and therefore $f+\iota_{T_{a}}=f$ and $g+\iota_{T_{a}}=g$.
For a proper convex function $f$, the singular subdifferential for some $x\in\mathbb{R}^{n}$ is defined as $\partial^{\infty}f(x):=N_{\operatorname{dom}(f)}(x)$ when $x\in\operatorname{dom}(f)$, and is empty when $x\notin\operatorname{dom}(f)$ (see, e.g. [10, 15]). Singular subdifferentials capture the notion of an infinite slope in a particular direction.
Theorem 5.1 (Qualification-free subdifferential additivity).
For $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ proper convex, $x\in\operatorname{dom}(f)\cap\operatorname{dom}(g)$, and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
$\partial(f+g)(x)=\partial(f+\iota_{T_{a}})(x)+\partial(g+\iota_{T_{a}})(x),$
$\partial^{\infty}(f+g)(x)=\partial^{\infty}(f+\iota_{T_{a}})(x)+\partial^{\infty}(g+\iota_{T_{a}})(x).$
Proof of Theorem 5.1.
By Corollary 4.3 applied to $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$,
$\operatorname{relint}\operatorname{dom}(f+\iota_{T_{a}})\cap\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})$ (3)
$=$ $\operatorname{relint}(\operatorname{dom}(f)\cap T_{a})\cap\operatorname{relint}(\operatorname{dom}(g)\cap T_{a})\neq\emptyset.$ (4)
Therefore, a qualification condition is satisfied for the functions $f+\iota_{T_{a}}$ and $g+\iota_{T_{a}}$, and the first part of the result follows by application of [14, Theorem 23.8].
The second part of the result follows from applying [14, Corollary 23.8.1] to the convex sets $\operatorname{dom}(f)\cap T_{a}$ and $\operatorname{dom}(g)\cap T_{a}$, by which
$N_{(\operatorname{dom}(f)\cap T_{a})\cap(\operatorname{dom}(g)\cap T_{a})}(x)$ $=N_{\operatorname{dom}(f)\cap T_{a}}(x)+N_{\operatorname{dom}(g)\cap T_{a}}(x)$
$=N_{\operatorname{dom}(f+\iota_{T_{a}})}(x)+N_{\operatorname{dom}(g+\iota_{T_{a}})}(x).$
And since $N_{(\operatorname{dom}(f)\cap T_{a})\cap(\operatorname{dom}(g)\cap T_{a})}(x)=N_{\operatorname{dom}(f)\cap\operatorname{dom}(g)}(x)=N_{\operatorname{dom}(f+g)}(x)$, the second part of the result follows. ∎
Theorem 5.2 (Normal cone of the intersection of convex sets).
For $C,D\subseteq\mathbb{R}^{n}$ convex, and $x\subseteq C\cap D$, for $T_{a}=T_{a}(C,D)$,
$N_{C\cap D}(x)=N_{C\cap T_{a}}(x)+N_{D\cap T_{a}}(x).$
Proof of Theorem 5.2.
With Corollary 4.3 we can use [14, Corollary 23.8.1], from which we find that
$N_{C\cap D}(x)=N_{(C\cap T_{a})\cap(D\cap T_{a})}(x)=N_{C\cap T_{a}}(x)+N_{D\cap T_{a}}(x).$
Recall that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the infimal convolution is
$(f\,\square g)(z):=\inf_{x,y\in\mathbb{R}^{n}:x+y=z}f(x)+g(y).$
In the following results, we denote by $f^{*}$ the convex conjugate of $f$.
Theorem 5.3 (Attained infimal convolution in the conjugate of a sum).
For two proper convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
$(f+g)^{*}=(f+\iota_{T_{a}})^{*}\square(g+\iota_{T_{a}})^{*},$
where the infimum in the infimal convolution is attained.
Proof of Theorem 5.3.
When $\operatorname{dom}(f)\cap\operatorname{dom}(g)=\emptyset$, $T_{a}=\emptyset$ and the equality holds trivially. Otherwise, by Corollary 4.3, $\operatorname{relint}(\operatorname{dom}(f+\iota_{T_{a}}))\cap\operatorname{relint}(\operatorname{dom}(g+\iota_{T_{a}}))\neq\emptyset$, and the result follows from [14, Theorem 20.1]. ∎
Theorem 5.4 (Generalized subdifferential chain rule).
Let $g\in\mathbb{R}^{m}\to\bar{\mathbb{R}}$ be a proper convex function, and $A\in\mathbb{R}^{m\times n}$ is a matrix. Then for $T=T(\operatorname{dom}(g),\operatorname{\mathrm{range}}(A))$,
$\partial(g)(Ax)=A^{T}\partial(g+\iota_{T})(Ax).$
Proof of Theorem 5.4.
Denote $T_{a}=T_{a}(\operatorname{dom}(g),\operatorname{\mathrm{range}}(A))$. Note that $\forall x\in\mathbb{R}^{n}$,
$(g\circ A)(x)=([g+\iota_{T_{a}}]\circ A)(x),$
because $T_{a}\supseteq\operatorname{dom}(g)\cap\operatorname{\mathrm{range}}(A)$. So it suffices to calculate the subdifferential of the r.h.s. of the above equation.
If $\operatorname{dom}(g)\cap\operatorname{\mathrm{range}}(A)=\emptyset,$ then the result holds trivially because $g(Ax)\equiv\infty$, and $(g+\iota_{T_{a}})(Ax)\equiv\infty$ as well. Otherwise, by Corollary 4.3 $\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})\cap\operatorname{\mathrm{range}}(A)\neq\emptyset$. Additionally, note that the only face of $\operatorname{\mathrm{range}}(A)$ is itself, whence by Proposition 4.1 $\operatorname{\mathrm{range}}(A)\subseteq T_{a}$. Since $0\in\operatorname{\mathrm{range}}(A)$, this in fact implies that $T_{a}=T$. The statement then follows by applying [14, Theorem 23.9] to the composition $g+\iota_{T}]\circ A$. ∎
In the following result, for a matrix $A\in\mathbb{R}^{m\times n}$ and a set $S\subseteq\mathbb{R}^{m}$, we denote by $A^{-1}(S)$ the preimage of $S$ under the linear map associated with the matrix $A$, making no assumption on the invertibility of the matrix $A$.
Theorem 5.5 (Exact Fenchel-Rockafellar dual).
For $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ proper, lower semi-continuous convex functions, and let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$. Then
$\inf_{x\in\mathbb{R}^{n}}f(x)+g(Ax)=-\inf_{y\in\mathbb{R}^{m}}[(f+\iota_{A^{-1}(T_{a})})^{*}(A^{T}y)+(g+\iota_{T_{a}})^{*}(-y)].$
Proof of Theorem 5.5.
In this proof we use the notation of left multiplication of functions by a matrix
$(Af)(u):=\inf\{f(x)\mid Ax=u\}.$
Note that $Af+\iota_{T_{a}}\equiv A(f+\iota_{A^{-1}(T_{a})})$. Indeed, both functions have the domain $(A\operatorname{dom}(f))\cap T_{a}$, where $A\operatorname{dom}(f):=\{Ax\mid x\in\operatorname{dom}(f)\}$, and $\forall y\in A\operatorname{dom}(f)\cap T_{a}$,
$(Af+\iota_{T_{a}})(y)$ $=\inf\{f(x)\mid Ax=y\}$
$=\inf\{f(x)+\iota_{Ax\in T_{a}}\mid Ax=y\}$
$=A(f+\iota_{A^{-1}(T_{a})})(y).$
Then, employing the change of variable $u=Ax$,
$\inf_{x\in\mathbb{R}^{n}}f(x)+g(Ax)$ $=\inf_{u\in\mathbb{R}^{m}}(Af)(u)+g(u)$
$=\inf_{u\in\mathbb{R}^{m}}(Af+\iota_{T_{a}})(u)+(g+\iota_{T_{a}})(u)$
$=\inf_{u\in\mathbb{R}^{m}}(A(f+\iota_{A^{-1}{T_{a}}}))(u)+(g+\iota_{T_{a}})(u)$
$=\inf_{x\in\mathbb{R}^{n}}(f+\iota_{A^{-1}{T_{a}}})(x)+(g+\iota_{T_{a}})(Ax).$
Note that by Corollary 4.3, $\exists u\in\operatorname{relint}\operatorname{dom}(Af)\cap\operatorname{relint}\operatorname{dom}(g)$. Note that $\operatorname{dom}(Af)=A\operatorname{dom}(f)$, and by [14, Theorem 6.6], $\operatorname{relint}A\operatorname{dom}(f)=A\operatorname{relint}\operatorname{dom}(f)$. Therefore, $u\in A\operatorname{relint}\operatorname{dom}(f)\cap\operatorname{relint}\operatorname{dom}(g)$, i.e. there exists an $x\in\operatorname{relint}\operatorname{dom}(f)$ such that $Ax\in\operatorname{relint}\operatorname{dom}(g)$. With this, we satisfy a sufficient requirement of [14, Corollary 31.2.1], which guarantees strong duality and completes the proof. ∎
6 Nested normals and facial subspaces
Inspired by the lexicographic characterization of convex faces of [8], we develop a framework to describe the lattice of convex faces at a point $x\in C$. By using this theory, we will then be able to show the results of Section 4.
We begin with a characterization of faces of convex sets. In the next few results, we denote $X$ for a real vector space.
Theorem 6.1 (Facial subspace characterisation of convex faces).
Let $C\subseteq X$. A subset $F\subseteq C$ is a face of $C$ if, and only if the two following conditions hold.
  1. 1.
    $\operatorname{aff}(F)\cap C=F$,
  2. 2.
    $C\setminus F$ is convex.
Proof of Theorem 6.1.
Let $F$ be a face. We argue by contradiction that $\operatorname{aff}(F)\cap[C\setminus F]=\emptyset$; let us assume that there is some $z\in\operatorname{aff}(F)\cap[C\setminus F]$.
If $F$ is a singleton, then $\operatorname{aff}(F)=F$, and $\operatorname{aff}(F)\cap[C\setminus F]=\emptyset$. Otherwise, we may assume that $0\in F\setminus\{z\}$ without loss of generality, because the faces of a convex set are invariant under translation. With this assumption, $\operatorname{aff}(F)=\operatorname{\mathrm{span}}(F)=\operatorname{cone}(F-F)$. Then $z\in\operatorname{cone}(F-F)$, so there are $y_{1},y_{2}\in F$ such that $z=\alpha(y_{1}-y_{2})$ for some $\alpha\geq 0$. We can rule out the possibility that $z\in[y_{1},y_{2}]$, because by convexity of $F$ it would follow that $z\in F$. The remaining possibility is that $y_{1}\in[y_{2},z[$. Further, note that $y_{1}\neq y_{2}$, because otherwise $z=0$, which disagrees with our assumption that $z\in F\setminus\{0\}$. Therefore, $y_{1}\in]y_{2},z[$. The line segment $y_{2},z]$ has a point $y_{1}\in]y_{2},z[\cap F$. Since $y_{2},z\in C$, the fact that $F$ is a face implies that $z$ must then be in $F$, in contradiction with our definition of $z$. We have therefore shown the claim.
We next argue that $C\setminus F$ is convex. $\forall y,z\in C\setminus F$, any $x\in]y,z[$ is in $C$ by convexity of $C$. Also, $x$ is not in $F$ because $F$ is a face. So any $x\in]y,z[$ is also in $C\setminus F$, which means that $C\setminus F$ is convex.
We now show the converse: let $F\subseteq C$ be such that $C\setminus F$ is convex and $\operatorname{aff}(F)\cap C=F$, we show that $F$ is a face of $C$.
Take any $y,z\in C$, $x\in F\cap]y,z[$. If $y\in\operatorname{aff}(F)$ then $z\in\operatorname{aff}(F)$ as well, because $z=\alpha(x-y)+y$ for some $\alpha>0$. So either $z,y$ are both in $\operatorname{aff}(F)$, or neither are. But if neither are, then $z,y\in C\setminus\operatorname{aff}(F)$, and by convexity of $C\setminus\operatorname{aff}(F)$, $x\in C\setminus\operatorname{aff}(F)$ also, which is a contradiction with $x\in F$. Therefore, $z,y$ are both in $F$, which shows that $F$ is a face. ∎
This characterization, surprisingly, appears to be novel to the knowledge of the author. It allows us to shift our focus away from the faces themselves to the comparatively simpler affine subspaces which select them through intersections with the convex set. We call such face-selecting subspaces affine facial subspaces.
Definition 6.1 (Affine facial subspace).
Let $C\subseteq X$ be convex. An affine subspace $L\subseteq V$ is an affine facial subspace of $C$ if $C\setminus L$ is convex.
The next result describes the relationship between affine facial subspaces and faces.
Corollary 6.1 (Facial subspace relation to faces).
Let $C\subseteq X$ be convex and $S\subseteq C$ non-empty.
  1. 1.
    For every affine facial subspace $L$ of $C$, the set $F:=C\cap L$ is a face of $C$.
  2. 2.
    Conversely, for every face $F$ of $C$, there exists an affine facial subspace $L$ such that $F=L\cap C$ (specifically, $L=\text{aff}(F)$).
Proof of Corollary 6.1.
For the first part of the result, we show that if $L$ is an affine facial subspace, then $F:=C\cap L$ satisfies the conditions of Theorem 6.1.
  1. 1.
    That $\operatorname{aff}(F)\cap C=F$: note that $\operatorname{aff}(F)\subseteq L$ because $F\subseteq L$, and so $\operatorname{aff}(F)\cap C\subseteq L\cap C=:F$. On the other hand, $L\cap C=(L\cap C)\cap C\subseteq\operatorname{aff}(L\cap C)\cap C$, and so $F\subseteq\operatorname{aff}(F)\cap C$.
  2. 2.
    That $C\setminus F$ is convex: note that
    $C\setminus F=C\setminus[C\cap L]=C\setminus L,$
    which is convex by definition of $L$ as an affine facial subspace.
We show the second part of the result. For $F$ a face of $C$, by Theorem 6.1 $F=C\cap\operatorname{aff}(F)$ and $C\setminus F$ is convex. But then $C\setminus\operatorname{aff}(F)=C\setminus(\operatorname{aff}(F)\cap C)=C\setminus F$, and therefore $C\setminus\operatorname{aff}(F)$ is convex, which makes $\operatorname{aff}(F)$ an affine facial subspace. ∎
In this work, we focus on “local” affine facial subspaces; the subset of affine facial subspaces that contain a point $x\in C$. These subsets have some additional nice algebraic properties. We find it useful to generalize this idea of “locality” to affine facial subspaces that are localized at a set $S\subseteq C$ instead of only at points, i.e., we consider all affine subspaces that contain $S$. To emphasize this connection, we use the word “at” to describe the containment of $S$: ”a facial subspace is at SS”. The reader should keep in mind the important special case of $S$ being a singleton $S=\{x\}$, for which the results are localized at a point.
Any affine facial subspace $L$ at $S$ has its affine offset determined by $S$. Recall that LLL-L is simply the affine subspace $L$ translated in such a way that it becomes a (non-affine) subspace. With this, it becomes clear that L=(LL)+SL=(L-L)+S. We can always deduce the correct affine offset from in $S$. Therefore, we find it convenient to describe the set of affine facial subspaces that contain $S$ through their translated, non-affine counterparts.
Definition 6.2 (Facial subspace at S).
Let $C\subseteq X$ be a convex set and $S\subseteq X$ a non-empty subset. A subspace $U\subseteq X$ is defined to be a facial subspace of $C$ at $S$ if:
  1. 1.
    $S-S\subseteq U$
  2. 2.
    $U+S$ is an affine facial subspace of $C$.
The set of all facial subspaces of $C$ at $S$ is denoted $\mathcal{F}(C;S)$.
When $S$ is the singleton $\{x\}$, the requirement that $S-S\subseteq U$ is trivially satisfied. When $S$ is not a singleton, this assumption guarantees that $\forall s_{1},s_{2}\in S,U+s_{1}=U+s_{2}$. This implies that $\forall s\in S,U+s=U+S$, and so the Minkowski sum $U+S$ is a simple affine offset of $U$, for which $S\subseteq U+S$. In other words, the affine facial subspace $U+S$ contains $S$.
Next, we introduce a generalization of the notion of a normal cone. A normal cone of a convex set $C\subseteq\mathbb{R}^{n}$ is defined as
$N_{C}(x):=\{z\in\mathbb{R}^{n}\mid\forall y\in C\langle y-x,z\rangle\leq 0\}.$
When $C$ is a cone and $x=0$, the normal cone $N_{C}(0)=:C^{\circ}$ is called the polar of $C$. The polar is a dual operation, in the sense that
${C^{\circ}}^{\circ}=\operatorname{cl}(C).$
We can see that a normal cone is a dual object which “describes” the original set only up to its closure.
We require a dual object which has sufficient descriptive power to go beyond the closure, encoding information about the boundary. To this end, we introduce the notion of a nested normal cone. Let $C\subseteq\mathbb{R}^{n}$ be convex and $U\subseteq\mathbb{R}^{n}$ be a subspace, then $\forall x\in\mathbb{R}^{n}$,
$N_{C}^{U}(x):=N_{C\cap U}(x)\cap U.$
We would also like for our normal cones to located at a set $S\subseteq C$, similarly to facial subspaces.
$N_{C}(S):=\{v\in\mathbb{R}^{n}\mid\forall s\in S,c\in C,\langle c-s,v\rangle\leq 0\}.$
The hyperplanes corresponding to the vectors in this normal cone are supporting hyperplanes of $C$ that contain $S$. Note also that $N_{C}(S)=\bigcap_{s\in S}N_{C}(s).$ The idea of normal cones that are located at sets appears in the earlier work [7], where normal cones are at faces of polytopes, instead at points. The normal cone at a set has an elegant connection to the notion of relative interior of that set: the normal cone at any point in the relative interior of a convex set $S\subseteq C$ will match the normal cone of $S$.
Combining the two generalizations, we get the following.
Definition 6.3 (nested normals).
Let $C\subseteq\mathbb{R}^{n}$ be a convex set, and $S$ a non-empty subset of $C$. Then for a subspace $U\subseteq\mathbb{R}^{n}$ such that $S-S\subseteq U$, and with $C^{\prime}:=C\cap(U+S)$,
$N_{C}^{U}(S):=\{v\in U\mid\forall c\in C^{\prime},s\in S,\langle c-s,v\rangle\leq 0\}$
is a nested normal cone of $C$ at $S$, and we label any $v\in N_{C}^{U}(S)$ to be a nested normal of $C$ at $S$ in $U$.
The idea of normal cone defined at a set has an implicit relation with the notion of relative interior.
Lemma 6.1 (nested normals in the relative interior are representative).
Let $C\subseteq\mathbb{R}^{n}$ be convex and $S\subseteq C$ a non-empty convex set. Let $U\subseteq\mathbb{R}^{n}$ be a subspace, and $x\in\operatorname{relint}(S)$. Then
$N_{C}^{U}(S)=N_{C}^{U}(x).$
Proof of Lemma 6.1.
That $N_{C}^{U}(S)\subseteq N_{C}^{U}(x)$ follows from simple inclusion $x\subseteq S$.
For the containment ($\supseteq$), WLOG let $U=\mathbb{R}^{n}$. We can make this assumption because the normal cones are invariant under the translation $U+S\to U$, and so we have an equivalent normal cone inside the subspace $U$. Since $U$ is a finite dimensional vector space, it is isomorphic to $\mathbb{R}^{n}$ for some $n\in\mathbb{N}$.
Let $v\in N_{C}(x)$. Then $\forall s\in S,\langle v,s\rangle\leq\langle v,x\rangle$ because $S\subseteq C$. We show that this inequality is in fact always an equality. Arguing by contradiction, assume that there is an $s\in S$ such that $\langle v,s\rangle<\langle v,x\rangle$. Then since $x\in\operatorname{relint}(S)$, there is a $\varepsilon>0$ such that $(\varepsilon B+x)\cap\operatorname{aff}(S)\subseteq S$ for $B$ the unit ball in $\mathbb{R}^{n}$. There is a $\lambda>0$ such that $s^{\prime}:=x+\lambda(s-x)\in\varepsilon B+x$. Then note that $s^{\prime}\in(\varepsilon B+x)\cap\operatorname{aff}(S)\subseteq S$, and
$\langle v,s^{\prime}\rangle=\langle v,x\rangle-\lambda\langle v,s-x\rangle.$
The second term is strictly positive, therefore $\langle v,s^{\prime}\rangle>\langle v,x\rangle$. But this is a contradiction with $v\in N_{C}(x)$. Therefore, we find that $\forall s\in S,$ $\langle v,s\rangle=\langle v,x\rangle$. Then $\forall c\in C,s\in S$ $\langle c-s,v\rangle=\langle c-x,v\rangle\leq 0$, and therefore $v\in N_{C}(S)$. ∎
The nested normals at a set are orthogonal with that set.
Lemma 6.2 (Matching sets means paired normals).
Let $C\subseteq\mathbb{R}^{n}$ be a convex set, and $U\subseteq\mathbb{R}^{n}$ a subspace. Then for any non-empty $S\subseteq C$ such that $S-S\subseteq U$,
$N_{C}^{U}(S)\subseteq(S-S)^{\perp}\cap U.$
Further, when $S=C\cap(U+S)$, the sets are equal.
Proof of Lemma 6.2.
Let $y\in N_{C}^{U}(S)$. By Definition 6.3, $\forall z,x\in S,$ since $x\in S$ and $z\in C$, $\langle x-z,y\rangle\leq 0,$ and since $x\in C$ and $z\in S$ $\langle x-z,y\rangle\geq 0$. Then $y\perp S-S$, which shows the inclusion.
Next we show the containment ($\supseteq$), in the case that $S=C\cap(U+S)$. Let $y\in(S-S)^{\perp}\cap U$. For any $x\in S,z\in C\cap(U+S)$ we need to show that $\langle z-x,y\rangle\leq 0$. Note that $z\in S$, so $x-z\in S-S$, and by definition of $y$, $\langle z-x,y\rangle=0$, which proves the statement. ∎
As previously discussed, one motivating property of the nested normals is that they can encode information about the boundary of convex sets. We make this idea concrete.
The tangent cone of a convex set $C$ at $x\in C$ is the smallest closed cone $T$ such that $C\subseteq x+T$. We can think of this cone as a local description of $C$ around $x$. The normal cone is the polar of hthe tangent cone, and therefore “describes” the convex set $C$ around $x$ to the same extent as the tangent cone. In contrast, nested normals in facial subspaces of $C$ at $x$ describe the cone of feasible directions.
Definition 6.4 (Cone of feasible directions).
For a convex set $C\in\mathbb{R}^{n}$ with $x\in C$, we define the cone of feasible directions of $C$ at $x$ to be $D_{C}(x):=\{v\in\mathbb{R}^{n}\mid\exists\lambda>0:\lambda v\in C\}$.
The closure of the cone of feasible directions is the tangent cone. The difference is that the cone of feasible directions captures additional information about the boundary of $C$ around $x$. The following result shows that any convex cone (including the cone of feasible directions at any point) can be recovered from the nested normals in facial subspaces.
Theorem 6.2 (Local polar without closure).
Given any convex cone $T$ with $0\in T$,
$T=\{x\in\mathbb{R}^{n}\mid\forall U\in\mathcal{F}(T;0),v\subseteq N_{T}^{U}(0),x\in U\implies\langle x,v\rangle\leq 0\}.$
Proof of Theorem 6.2.
Let $T^{\prime}:=\{x\in\mathbb{R}^{n}\mid\forall U\in\mathcal{F}(T;0),v\subseteq N_{T}^{U}(0),x\in U\implies\langle x,v\rangle\leq 0\}$. By the definition of conditional normals, $T\subseteq T^{\prime}$. For the containment ($\supseteq$), we argue by contradiction: we assume that there is some $x\in T^{\prime}\setminus T$. If $x\in\mathrm{int}(T^{c})$, then $x$ is strictly separated from $T$ by a supporting hyperplane of $T$ (e.g., see [14, Theorem 11.3]). Any such hyperplane induces a nested normal of the form $v\in N_{T}^{\mathbb{R}^{n}}(0)$, for which $v^{\perp}$ is the hyperplane, and $\langle v,x\rangle>0$. But then $x\notin T^{\prime}$, and so $x\notin\mathrm{int}(T^{c})$. Since we assumed that $x\notin T$, the only possibility is that $x\in\partial T$. Then $x$ is contained in a supporting hyperplane, so there is some nested normal $v\in N_{T}^{\mathbb{R}^{n}}(0)$ such that $\langle v,x\rangle=0$. Let $T_{1}=T\cap v^{\perp}$, and narrow the ambient space to the subspace $T_{1}$. Inside $T_{1}$, we repeat the same argument, and keep narrowing the ambient space in this manner. This process must eventually reduce the ambient space to $\operatorname{\mathrm{span}}(x)$, because $\mathbb{R}^{n}$ is finite-dimensional. In this one-dimensional space, with the narrowed cone $T_{\ell}=T\cap\operatorname{\mathrm{span}}(x)$, note that $T_{\ell}$ is a closed cone. Therefore, either $x\in T_{\ell}$ or $x\in\mathrm{int}(T_{\ell}^{c})$, both of which yields a contradiction, from which the result follows. ∎
Another important property of nested normals is that they describe the containment structure of facial subspaces, which is closely related to the lattice structure of the faces. We make this precise in the next two results.
Lemma 6.3 (Nested normals induce facial subspaces).
Let $C\subseteq\mathbb{R}^{n}$ be convex, $S\subseteq C$ non-empty, $U\subseteq\mathbb{R}^{n}$ a subspace, and $v\in N_{C}^{U}(S)$. Let $V\in\mathcal{F}(C;S)$ be such that $V\subseteq U$. Then
$V\cap v^{\perp}\in\mathcal{F}(C;S).$
Lemma 6.4 (Nested normals induce facial subspaces).
First, note that if $v=0$, then $V\cap v^{\perp}=V$, and there result holds. We now consider the case where $v\neq 0$.
By Lemma 6.2, $v\perp S-S$, and therefore $S-S\subseteq V\cap v^{\perp}$.
We make two observations.
First, since $v\in N_{C}^{U}(S)$ and $V\subseteq U$, $\forall y\in C\cap(V+S)$, $\forall s\in S,$ we have that $\langle y,v\rangle\leq\langle s,v\rangle$. Second, this inequality is tight precisely for those values of $y$ that are also contained in $v^{\perp}+S$, i.e.,
$\{y\in C\cap(V+S)\mid\forall s\in S,\langle y,v\rangle=\langle s,v\rangle\}$ (5)
$=$ $((V+S)\cap C)\cap(v^{\perp}+S)=(V\cap v^{\perp}+S)\cap C.$ (6)
These two facts imply that
$(V\cap v^{\perp}+S)\cap C=\operatorname*{\arg\max}_{y\in C\cap(V+S)}\langle y,v\rangle.$
The r.h.s. is an exposed face of the face $C\cap(V+S)$, so it is itself a face, i.e., $V\cap v^{\perp}\in\mathcal{F}(C;S)$.
We next show that all facial subspaces are “reached” by nested normals in facial subspaces.
Lemma 6.5 (Completeness of conditional normal proof composition).
Let $C\subseteq\mathbb{R}^{n}$ be convex, $S\subseteq C$ non-empty, and $V,U\in\mathcal{F}(C;S)$ such that $U\subseteq V$. Then there is a sequence $V_{0},\ldots,V_{k}\in\mathcal{F}(C;S)$ with $k=\dim(V)-\dim(U)$ such that $V_{0}=V$, $V_{k}=U$, and for each $i\in\{1,\ldots,k\}$ there is a $v_{i}\in N_{C}^{V_{i-1}}(S)\setminus\{0\}$ such that $V_{i}=V_{i-1}\cap v_{i}^{\perp}$.
This result is similar to the lexicographic characterization of the faces of convex sets in [8, Proposition 6], which states that every face can be obtained by a sequence of faces where each face is an exposed face of the preceding face. Let us state a technical lemma that we will need for the proof.
Lemma 6.6 (Affine slices preserve facial subspaces).
Let $C\subseteq\mathbb{R}^{n}$ be convex and $S\subseteq C$ non-empty. Let $U\in\mathcal{F}(C;S)$, and $V$ any subspace such that $S-S\subseteq V$. Then $U\cap V\in\mathcal{F}(C\cap(V+S);S)$.
Proof of Lemma 6.6.
Since $C\setminus(U+S)$ is convex, $(V+S)\cap(C\setminus(U+S))$ is also convex. We can write this set as $(V+S)\cap C]\setminus[(V+S)\cap(U+S)]$, which is the same as $((V+S)\cap C)\setminus((V\cap U)+S)$, because $S-S\in V\cap U$. Since this set is convex, and $S-S\subseteq V\cap U$, we find that $V\cap U\in\mathcal{F}((V+S)\cap C;S)$. ∎
Proof of Lemma 6.5.
It suffices to show the following statement: given any $U,V\in\mathcal{F}(C;S)$ such that $V\supsetneq U$, there exists a $v\in N_{C}^{V}(S)\setminus\{0\}$ such that $v\perp U$. If this statement is shown to hold, applying it recursively generates the sequence of facial subspaces $V_{1},\ldots,V_{k}$ in a finite number of steps. Note that the dimension of the facial subspaces decreases by one at each iteration while always containing $U$, which implies that $U\subseteq V_{k}$ with $\dim(U)=\dim(V_{k})$, and therefore $U=V_{k}.$
We now show the above statement. Let $C^{\prime}=C\cap(V+S)$, $\bar{C}=\operatorname{\Pi}_{U^{\perp}}C^{\prime}$, and $\bar{U}=\operatorname{\Pi}_{U^{\perp}}(U+S)$. Note that $\bar{C}$ is a convex set and $\bar{U}$ a singleton $\{x\}\subseteq\bar{C}$. We argue that $\bar{U}$ is a face of $\bar{C}$. Arguing by contradiction, we hypothesize the $\bar{U}$ is not a face of $\bar{C}$, in which case there are $\bar{y},\bar{z}\in\bar{C}\setminus\bar{U}$ such that $x=\lambda\bar{y}+(1-\lambda)\bar{z}$ for $\lambda\in(0,1)$. This implies that there are elements $y,z\in C^{\prime}$ such that $\bar{y}=\operatorname{\Pi}_{U^{\perp}}y$ and $\bar{z}=\operatorname{\Pi}_{U^{\perp}}z$, with $w=\lambda y+(1-\lambda)z$. Note that the line segment $[y,z]$ has endpoints $y,z\in C^{\prime}\setminus(U+S)$ with a point $w\in]y,z[$ that is also contained in $C^{\prime}\cap(U+S)$. This implies that $C^{\prime}\cap(U+S)$ is not a face of $C^{\prime}$, which, by Corollary 6.1, means that $U\notin\mathcal{F}(C^{\prime};S)$. This is a contradiction because in fact, $U\in\mathcal{F}(C^{\prime};S)$, as we now show. We know that $U\in\mathcal{F}(C;S)$, so by Lemma 6.6, $U=U\cap V\in\mathcal{F}(C\cap(V+S);S)$, i.e., $U\in\mathcal{F}(C^{\prime};S)$. Therefore, we indeed have a contradiction, from which we find that $\bar{U}$ is a face of $\bar{C}$. Note that $\operatorname{\Pi}_{U^{\perp}}(V+S)\subseteq V+S$. This becomes apparent from the following equality: $V+S=(V+S)+U=\operatorname{\Pi}_{U^{\perp}}(V+S)+U,$ where the first equality holds because $U\subseteq V$ implies that $V=V+U$.
Then since $U+S,C\cap(V+S)\subseteq V+S$, we find that $\bar{U},\bar{C}\subseteq(V+S)\cap U^{\perp}$. Since $\bar{U}$ is a face of $\bar{C}$, there exists a supporting hyperplane $H$ of $\bar{C}$, where the hyperplane is relative to the affine subspace $(V+S)\cap U^{\perp}$, such that $\bar{U}\subseteq H$. Indeed, if there is no such hyperplane $H$, $\bar{U}$ would then be a singleton in the interior of $\bar{C}$ (where the interior is relative to $(V+S)\cap U^{\perp}$, i.e., with the induced topology in $(V+S)\cap U^{\perp}$).
Take $\bar{v}\in V\cap U^{\perp}$ to be the unit normal to $H$ that is also in $N_{\bar{C}}^{V\cap U^{\perp}}(x)$. Then note that $\bar{v}\in N_{C}^{V}(S)\cap U^{\perp}$, which is the vector we had to find. ∎
Together, Lemma 6.3 and Lemma 6.5 enables a powerful tool that is central to our proofs: we can perform structural induction on the tree of facial subspaces, using only nested normals. Indeed, the set of facial subspaces forms a tree, where the root is $\mathbb{R}^{n}$ (note that it is always a facial subpace, and that it contains any other facial subspace), and every edge of the tree is associated with a nested normal. The induction statement is the following.
Proposition 6.1 (Induction by nested normal).
Let $C$ be a convex set, $S\subseteq C$ a non-empty subset, and a statement $A:\mathcal{F}(C;S)\to\{\text{true},\text{false}\}$. If we have show that:
  1. 1.
    $A(\mathbb{R}^{n}).$
  2. 2.
    $\forall U\in\mathcal{F}(C;S)$, $v\in N^{U}_{C}(S)$, $A(U)\implies A(U\cap v^{\perp})$. Then it follows that
    $\forall V\in\mathcal{F}(C;S),\,A(W).$
6.1 Minimal facial subspace
Recall that the smallest face of a convex set $C$ containing some $x\in C$ is $F_{C}(x)$, the face of $C$ generated by $x$. It is the only face with $x$ in its relative interior. In this section we develop an analogous notion for facial subspaces, and characterize this facial subspace by using nested normals.
Definition 6.5 (The generated facial subspace).
Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, we define the generated facial subspace $H_{C}(S)$ of $C$, generated by $S\subseteq\mathbb{R}^{n}$, to be the facial subspace that is contained in every other facial subspace in $\mathcal{F}(C;S)$.
We show existence and uniqueness for the above definition, starting with a lemma.
Lemma 6.7 (Facial subspaces are closed under intersections).
For $C\subseteq\mathbb{R}^{n}$ convex and $S\subseteq C$ non-empty, let $U,V\in\mathcal{F}(C;S)$. Then $U\cap V\in\mathcal{F}(C;S)$.
Proof of Lemma 6.7.
The subspace $U\cap V$ is such that
$(U\cap V+S)\cap C=((U+S)\cap C)\cap((V+S)\cap C).$
The r.h.s. is the intersection of two convex faces of $C$. Their intersection is then also a face (this is immediate from Definition 3.1). By Theorem 6.1, $C\setminus[(U\cap V+S)\cap C]$ is convex. The convexity of this set, which we can write as $C\setminus((U\cap V)+S)$, implies that $U\cap V\in\mathcal{F}(C;S)$. ∎
Theorem 6.3 (Existence and uniqueness of the generated facial subspace).
The generated facial subspace of Definition 6.5 exists and is unique.
Proof of Theorem 6.3.
We argue that there exists a facial subspace which contains no other. Arguing for the existence by contradiction, if every facial subspace of $C$ at $S$ properly contains some other facial subspace of $C$ at $S$, starting at any facial subspace, we can find an infinite chain of facial subspaces with dimension decreasing by one at each step. This is not possible because the dimension is finite. Therefore this chain of subspaces must terminate at some subspace which contains no other.
We now argue that there is only one subspace that contains no other. Let $V_{1},V_{2}$ be two distinct minimal facial subspaces of $C$ at $S$. Then by Lemma 6.7, $V_{1}\cap V_{2}$ is a facial subspace of $C$ at $S$ properly contained in both $V_{1}$ and $V_{2}$, which contradicts $V_{1}$ and $V_{2}$ being minimal.
We have shown the existence and uniqueness of a facial subspace that contains no other. From any facial subspace, we can start a proper containment chain as in our existence argument. The subspace at the end of this chain is the subspace which contains no other, which concludes the proof. ∎
As might be expected, the generated facial subspace is closely related to the generated face.
Lemma 6.8 (Relation of the generated face to the generated facial subspace).
With $C\subseteq\mathbb{R}^{n}$ convex and $S\subseteq C$ non-empty,
  1. 1.
    $F_{C}(S)=C\cap(H_{C}(S)+S).$
  2. 2.
    $H_{C}(S)=\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S)).$
Proof of Lemma 6.8.
By Theorem 6.1, $H_{C}(S)$ being the smallest facial subspace implies that $C\cap(H_{C}(S)+S)$ is the smallest face containing $S$, because any properly contained, smaller face would have its own facial subspace (by Corollary 6.1), which would have to be properly contained in $H_{C}(S)$.
For the second statement, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is a facial subspace because $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))+S=\operatorname{aff}(F_{C}(S))$, which is an affine facial subspace, as implied by Theorem 6.1. Further, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is contained in any other facial subspace in $\mathcal{F}(C;S)$, because given $U\in\mathcal{F}(C;S)$ such that $U\subsetneq\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$, we would then have $C\cap(U+S)$ as a face of $C$ containing $S$ that is properly contained in $F_{C}(S)$, in contradiction with the definition of $F_{C}(S)$. ∎
We can now reformulate Definition 4.1 into an equivalent definition.
Definition 6.6 (Equivalent definition of the joint facial subspace).
The joint facial subspace of two convex sets $C,D\subseteq\mathbb{R}^{n}$ from Definition 4.1 is
$T(C,D):=H_{C-D}(0).$
The next result characterizes generated facial subspaces in terms of nested normals.
Lemma 6.9 (The generated facial subspace is the only facial subspace with no nested normals).
Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, the generated facial subspace $H_{C}(S)$ is the only facial subspace of $C$ at $S$ such that $N_{C}^{H_{C}(S)}(S)=\{0\}$.
Proof of Lemma 6.9.
The generated facial subspace does not contain any non-trivial nested normals, because any non-trivial nested normal $v\in N_{C}^{H(C)}(S)\setminus\{0\}$ induces the facial subspace $H_{C}(S)\cap v^{\perp}\subsetneq H_{C}(S)$ by Lemma 6.3.
For any facial subspace $V\in\mathcal{F}(C;S)$ such that $N_{C}^{V}(S)=\{0\}$, we argue that $V=H_{C}(S)$. Note that $H_{C}(S)\subseteq V$ by Definition 6.5. The inclusion cannot be proper, because otherwise, by Lemma 6.5, there would be a nontrivial nested normal $v\in N_{C}^{V}(S)\setminus\{0\}$. ∎
6.2 Nested normals of a difference of sets
In this section, we show the following important result.
Lemma 6.10 (Characterization of the conditional normals of a difference of sets).
Let $C,D\subseteq\mathbb{R}^{n}$ be convex with $S\subseteq C\cap D$ non-empty, and $U\in\mathcal{F}(C-D;0)$. Then $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$, and
$N^{U}_{C-D}(0)=N^{U}_{C}(S)\cap(-N^{U}_{D}(S)).$
To show Lemma 6.10, we require the following lemma.
Lemma 6.11 (Normal cone of the difference of convex sets).
Given convex sets $C,D\subseteq\mathbb{R}^{n}$ with $S\subseteq C\cap D$ non-empty,
$N_{C-D}(0)=N_{C}(S)\cap(-N_{D}(S)).$
Proof of Lemma 6.11.
We start with the inclusion $(\subseteq)$. Any $v\in N_{C-D}(0)$ has the property that $\forall c\in C,d\in D$, $\langle v,(c-d)-0\rangle\leq 0$. Then $\forall c\in C,s\in S$, $\langle v,c-s\rangle\leq 0$ because $s\in D$, and therefore $v\in N_{C}(S)$. By a similar argument, $v\in N_{D}(S)$, which shows the inclusion.
For the containment ($\supseteq$), fix some $s\in S$. let $v\in N_{C}(S)\cap(-N_{D}(S))$. Then $\forall c\in C,d\in D$
$\langle v,c-d\rangle=\langle v,c-s\rangle+\langle v,s-d\rangle\leq 0,$
and so $v\in N_{C-D}(0)$. ∎
Proof of Lemma 6.10.
We argue by structured induction over the facial subspaces of CDC-D, as described in Proposition 6.1. Fixing some $S\subseteq C\cap D$, the statement $A:\mathcal{F}(C-D;0)\to\{\text{true},\text{false}\}$ that we show is the following: $A(U)$ is true if, and only if, all of the following hold.
  1. 1.
    $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S).$
  2. 2.
    $N^{U}_{C-D}(0)=N^{U}_{C}(S)\cap(-N^{U}_{D}(S)).$
  3. 3.
    $C\cap(U+S)-D\cap(U+S)=(C-D)\cap U.$ The additional point 3) has a supportive role in this proof.
The base case, where $U=\mathbb{R}^{n}$, holds because $\mathbb{R}^{n}$ is always a facial subspace. Point 2) is given by Lemma 6.11, and point 3) holds trivially.
For the inductive step, let $V\in\mathcal{F}(C-D;0)$ be such that $A(V)$ is true. Let $v\in N_{C-D}^{V}(0)\setminus\{0\}$, and $U=V\cap v^{\perp}$. In the rest of the proof, we show that $A(U)$ is true. Note that by point 2) of $A(V)$, $v\in N_{C}^{V}(S)\cap(-N_{D}^{V}(S))$, which by Lemma 6.3 implies that $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$, and this shows point 1) of $A(U)$.
We next show that for $C^{\prime}=C\cap(V+S)$, $D^{\prime}=D\cap(V+S)$,
$(C^{\prime}-D^{\prime})\cap U=(C^{\prime}\cap(U+S))-(D^{\prime}\cap(U+S)).$ (8)
We start with the inclusion ($\subseteq$): Take $z\in C^{\prime}$, $y\in D^{\prime}$ with $z-y\in U$. Then since $U\subseteq v^{\perp}$, $\langle z-y,v\rangle=0$, and therefore $\forall s\in S$, $\langle z-s,v\rangle-\langle y-s,v\rangle=0$. That $\langle z-s,v\rangle\leq 0$ follows from $v\in N_{C}^{V}(S)$ since $z\in C^{\prime}\subseteq C$ and $z-s\in V$, and similarly, $\langle y-s,v\rangle\geq 0$ because $v\in N_{D}^{V}(S)$. But then the only way that $\langle z-s,v\rangle-\langle y-s,v\rangle=0$ is if $\langle z-s,v\rangle=0$ and $\langle y-s,v\rangle=0$ individually. It follows that $z,y\in U+S$, and so the inclusion holds. The containment ($\supseteq$) is trivial, so we have shown Equation 8.
By point 3) of $A(V)$, $C^{\prime}-D^{\prime}=(C-D)\cap V$. Also, since $U\subseteq V$, $C^{\prime}\cap(U+S)$ = $C\cap(U+S)$, and similarly for $D$. Applying these identities on both sides of Equation 8, we find that
$(C-D)\cap U=C\cap(U+S)-D\cap(U+S).$
This shows that point 3) of $A(U)$ is true.
To show point 2) of $A(U)$, we compute
$N_{C}^{U}(S)\cap(-N_{D}^{U}(S))$ $=[N_{C\cap(U+S)}(S)\cap(-N_{D\cap(U+S)}(S))]\cap U$ (9)
$=N_{C\cap(U+S)-D\cap(U+S)}(0)\cap U$ (10)
$=N_{(C-D)\cap U}(0)\cap U$ (11)
$=N_{C-D}^{U}(0),$ (12)
where Equation 9 holds by definition of nested normals, Equation 10 follows from Lemma 6.11 applied to the sets $C\cap(U+S)$ and $D\cap(U+S)$, and Equation 11 follows from point 3) of $A(U)$. This shows that point 2) of $A(U)$ holds for $U$, and the proof is complete. ∎
We now have the tools to prove Corollary 4.1.
Proof of Corollary 4.1.
Considering Definition 4.1 in the light of Corollary 6.1, $T(C,D)=H_{C-D}(0)$. So $T\in\mathcal{F}_{C-D}(0)$, which by Lemma 6.10 implies that, taking $S=C\cap D$, $T\in\mathcal{F}_{C}(C\cap D)\cap\mathcal{F}_{D}(C\cap D)$. By the definition of facial subspaces (Definition 6.2), $C\cap D-C\cap D\subseteq T$. ∎
7 Proofs
We first show a smaller result, after which each subsection is dedicated to proving one result from Section 4.
Proof of Proposition 5.1.
First we show that for $C\subseteq\mathbb{R}^{n}$ convex, and $x\in\operatorname{relint}(C)$, $F_{C}(x)=C$. Note that
$N_{C}^{\operatorname{\mathrm{span}}(C-C)}(x)=N_{C}^{\operatorname{\mathrm{span}}(C-C)}(C)=(C-C)^{\perp}\cap\operatorname{\mathrm{span}}(C-C)=\{0\},$
where the first equality follows by Lemma 6.1 and the second by Lemma 6.2. Then by Lemma 6.9, $H_{C}(x)=\operatorname{\mathrm{span}}(C-C)$, and by Lemma 6.8, $F_{C}(x)=(\operatorname{\mathrm{span}}(C-C)+x)\cap C=C$.
From Proposition 4.1, for $x\in\operatorname{relint}(\operatorname{dom}(f))\cap\operatorname{relint}(\operatorname{dom}(g))$ and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
$T_{a}\cap\operatorname{dom}(f)=F_{\operatorname{dom}(f)}(\operatorname{dom}(f)\cap\operatorname{dom}(g))\supseteq F_{\operatorname{dom}(f)}(x)=\operatorname{dom}(f).$
From this the results follows. ∎
7.1 Regularity in the joint facial subspace
We now show the main property of the joint facial subspace, Theorem 4.2. We first state and prove a lemma which connects the notion of separation of convex set to a statement reminiscent of the qualification condition by Mordukhovich in variational analysis: [10, Theorem 2.19].
Lemma 7.1 (Equivalence of collisions and separation).
For convex sets $C,D\subseteq\mathbb{R}^{n}$, and $S\subseteq C\cap D$ non-empty, the following statement holds.
$N_{C}(S)\cap(-N_{D}(S))\neq\{0\}\iff C\text{ is separated from }D.$ (13)
Proof of Lemma 7.1.
We start with $(\implies)$. There exists a $v\in[N_{C}(S)\cap(-N_{D}(S))]\setminus\{0\}$, and from Lemma 6.2 $v\perp S-S$, whence the hyperplane $v^{\perp}+S$ separates $C$ from $D$.
[…]
  1. 1.
    $\forall s\in S,\langle v,s\rangle=\beta$.
  2. 2.
    $\forall c\in C$, $\langle y,v\rangle\geq\beta$.
  3. 3.
    $\forall d\in D$, $\langle d,v\rangle\leq\beta$. Then $\forall y\in C,s\in S$, $\langle v,y-s\rangle\geq 0,$ and so $v\in N_{C}(S)$. Similarly, $v\in N_{D}(S)$, so $v\in N_{C}(S)\cap(-N_{D}(S))$, which shows that $N_{C}(S)\cap(-N_{D}(S))\neq\{0\}$.
Proof of Theorem 4.2.
Recall that from Definition 6.6, $T=H_{C-D}(0)$. Then by Lemma 6.9, $N_{C-D}^{T}(C\cap D)=\{0\}$. Applying Lemma 6.10, we find that $N^{T}_{C}(C\cap D)\cap(-N^{T}_{D}(C\cap D))=\{0\}$.
By applying Lemma 7.1 in the space $T_{a}$, the result follows. Note that Lemma 7.1 is stated in $\mathbb{R}^{n}$, yet we can apply it in $T_{a}$. Indeed, $T$ is a finite-dimensional subspace, and thereby isomorphic to $\mathbb{R}^{n}$, and Lemma 7.1 is invariant when translating $T$ to $T_{a}$. ∎
7.2 First characterization
In this section we prove Proposition 4.1 and Theorem 4.1.
Lemma 7.2 (nested constraint subspaces for nested normals).
Let $C\subseteq\mathbb{R}^{n}$ be a convex set, $S\subseteq C$ be non-empty, and $U,V\subseteq\mathbb{R}^{n}$ be subspaces such that $S-S\subseteq U\subseteq V$ and $C\cap(V+S)\subseteq U+S$. Then
$N_{C}^{U}(S)+(U^{\perp}\cap V)=N_{C\cap(U+S)}^{V}(S).$
Proof of Lemma 7.2.
We start with the containment ($\supseteq$). Let $x\in N_{C}^{V}(S).$ Then $\forall s\in S,y\in C\cap(V+S)$, note that $y-s\in U$, and so
$\langle\operatorname{\Pi}_{U}x,y-s\rangle=\langle x,y-s\rangle\leq 0.$
This shows that $\operatorname{\Pi}_{U}x\in N_{C}^{U}(S)$. Now $x=\operatorname{\Pi}_{U}x+\operatorname{\Pi}_{U^{\perp}}x$, and $x,\operatorname{\Pi}_{U}x\in V\implies\operatorname{\Pi}_{U^{\perp}}x\in V$, therefore $x\in N_{C}^{U}(S)+(U^{\perp}\cap V)$.
For the inclusion ($\subseteq$), we consider any $q\in N^{U}_{C}(S)$ and $z\in U^{\perp}\cap V$. Let $y\in C\cap(U+S),s\in S$, and note that $\langle z,y-s\rangle=0$ because $y-s\in U$. Then
$\langle q+z,y-s\rangle=\langle q,y-s\rangle\leq 0.$
Therefore $q+z\in N_{C\cap(U+S)}^{V}(S)$, which shows the inclusion and completes the proof. ∎
Lemma 7.3 (Conical projections yield facial subspaces).
Let $C\subseteq\mathbb{R}^{n}$ be convex, $S\subseteq C$ non-empty, and $U\in\mathcal{F}(C;S)$. Then for any $K\subseteq N^{U}_{C}(S)$,
$U\cap K^{\perp}\in\mathcal{F}(C;S).$
Proof of Lemma 7.3.
Since $K$ is finite dimensional, there is a finite number of vectors $v_{1},\ldots,v_{m}\in K$ such that $\operatorname{\mathrm{span}}(K)=\operatorname{\mathrm{span}}(v_{1},\ldots,v_{m})$. Then since $v_{1},\ldots,v_{m}\in N_{C}^{U}(S)$, by repeated application of Lemma 6.3, $U\cap v_{1}^{\perp}\ldots\cap v_{m}^{\perp}$ is a facial subspace, which is the same set as $U\cap K^{\perp}$. ∎
Lemma 7.4 (Normal to the generated facial subspace of one convex set).
For convex sets $C,D\subseteq\mathbb{R}^{n}$, $T=T(C,D)$, $T_{a}=T+C\cap D$, we have
$H_{C}(C\cap D)=T\cap N_{C}^{T}(C\cap T_{a})^{\perp}.$
Proof of Lemma 7.4.
Let $U=T\cap N_{C}^{T}(C\cap T_{a})^{\perp}$. we first show that $U\in\mathcal{F}(C\cap D;C)$. Note that $T\in\mathcal{F}(C\cap D;C)$ by Lemma 6.10. Then we can use Lemma 7.3 because $N_{C}^{T}(C\cap T_{a})\subseteq N_{C}^{T}(C\cap D)$, which shows that $U\in\mathcal{F}(C\cap D;C)$ as well.
In the remainder of the proof we show that the assumption of Lemma 6.9 holds for $U$, namely that $N_{C}^{U}(C\cap D)=\{0\}$.
We first show that $N_{C}^{U}(C\cap T_{a})=\{0\}$. By Lemma 6.2, $N_{C}^{T}(C\cap T_{a})=(C\cap T_{a}-C\cap T_{a})^{\perp}\cap T$. Also, note that by the identity in Lemma 7.2, we find that $N_{C}^{U}(C\cap T_{a})=N_{C}^{T}(C\cap T_{a})\cap U$. Combining these two facts,
$N_{C}^{U}(C\cap T_{a})=N_{C}^{T}(C\cap T_{a})\cap U=(C\cap T_{a}-C\cap T_{a})^{\perp}\cap U.$
But note that $C\cap T_{a}-C\cap T_{a}\subseteq U$, therefore it fo $N_{C}^{U}(C\cap T_{a})=\{0\}$. That $C\cap T_{a}-C\cap T_{a}\subseteq U$ holds because $C\cap T_{a}-C\cap T_{a}\subseteq T$ and $(C\cap T_{a}-C\cap T_{a})\perp N_{C}^{T}(C\cap T_{a})$ by Lemma 6.2.
It only remains to show that $N_{C}^{U}(C\cap D)=N_{C}^{U}(C\cap T_{a})$. By Corollary 4.3, we may pick some $x\in\operatorname{relint}(C\cap T_{a})\cap\operatorname{relint}(D\cap T_{a})$. Then by Lemma 6.1,
$N_{C}^{U}(C\cap T_{a})=N_{C}^{U}(x)\supseteq N_{C}^{U}(C\cap D).$
Then $N_{C}^{U}(C\cap D)=\{0\}$, and so the result follows from Lemma 6.9. ∎
Proof of Proposition 4.1.
From Lemma 7.4, $H_{C}(C\cap D)=T\cap N_{C}^{T}(C\cap T_{a})^{\perp}$. Since $(C\cap T_{a}-C\cap T_{a})\perp N_{C}^{T}(C\cap T_{a})$ by Lemma 6.2, $(C\cap T_{a}-C\cap T_{a})\subseteq H_{C}(C\cap D)$, i.e., $H_{C}(C\cap D)+C\cap T_{a}\supseteq C\cap T_{a}$. But $H_{C}(C\cap D)+C\cap T_{a}=H_{C}(C\cap D)+C\cap D$, therefore
$H_{C}(C\cap D)+C\cap D\supseteq C\cap T_{a}.$ (14)
Additionally, since we know that $T\in\mathcal{F}(C\cap D;C)$ from Lemma 6.10, by definition of the generated facial subspace we know that $H_{C}(C\cap D)\subseteq T$. Therefore,
$H_{C}(C\cap D)+C\cap D\subseteq T_{a}.$ (15)
From Equation 14 and Equation 15 it follows that $(H_{C}(C\cap D)+C\cap D)\cap C=C\cap T_{a}$. Then the statement follows from Lemma 6.8. ∎
Proof of Theorem 4.1.
From Proposition 4.1, we have that $T\in\mathcal{F}(C;C\cap D)$ is a facial subspace of $C$ that contains $C\cap D$, and so by Definition 6.5, $H_{C}(C\cap D)\subseteq T$. Therefore, for $^{\prime}:=H_{C}(C\cap D)+H_{D}(C\cap D), TT$$^{\prime}\subseteq T$.
For the other containment, we argue by contradiction: let $T^{\prime}\subsetneq T$. Then there is a unit vector $v\in T$ with $v\perp T^{\prime}$. Then $v\perp H_{C}(C\cap D)$, which by Lemma 7.4, implies that $v\in N_{C}^{T}(C\cap T)$. By Lemma 6.2, $v\in N_{C}^{T}(C\cap T)$ as well. And since $N_{C}^{T}(C\cap T)\subseteq N_{C}^{T}(C\cap D)$, we find that $v,-v\in N_{C}^{T}(C\cap D)$. By a similar argument, $v,-v\in N_{D}^{T}(C\cap D)$. But then $v\in N_{C}^{T}(C\cap D)\cap(-N_{D}^{T}(C\cap D))$, which by Lemma 6.10 means that $v\in N_{C-D}^{T}(0)$. But by Lemma 6.9, $T$ cannot contain such a nested normal because by Definition 6.6, $T=H_{C-D}(0)$. Therefore, by contradiction $T=T^{\prime}$. ∎
7.3 Second characterization
We can now state the second characterization in full.
Theorem 7.1 (Iterative bilateral facial reduction).
Let $C,D\subseteq\mathbb{R}^{n}$ be convex sets and $S\subseteq C\cap D$ non-empty. Beginning with $T_{0}=\mathbb{R}^{n}$, let
$T_{i+1}=T_{i}\cap(N^{T_{i}}_{C}(S)\cap[-N^{T_{i}}_{D}(S)])^{\perp}.$
Let $\ell$ be the first index for which $T_{\ell+1}=T_{\ell}$. Then:
  1. 1.
    $T_{\ell}=T(C,D)$.
  2. 2.
    $\ell\leq n$.
Proof of Theorem 7.1.
First, $(N^{T_{i}}_{C}(S)\cap[-N^{T_{i}}_{D}(S)])=N^{T_{i}}_{C-D}(S)$ by Lemma 6.10. Applying Lemma 7.3 at each step, we find that $\forall i\in\{0,\ldots,\ell\},\,T_{i}\in\mathcal{F}(C-D;S)$. Finally, note that whenever $N^{T_{i}}_{C}(S)\cap[-N^{T_{i}}_{D}(S)]\neq\{0\}$, $\dim(T_{i+1})\leq\dim(T_{i})-1$. Therefore, the dimension may be reduced at most $n$ times, which implies that $\ell\leq n$. On the other hand, the fact that $T_{\ell}=T_{\ell+1}$ implies that $N^{T_{\ell}}_{C-D}(S)=\{0\}$. Then from Lemma 6.9 it follows that $T_{\ell}=H_{C-D}(S)$, which is $T(C,D)$ from Definition 6.6. ∎
7.4 Aknowledgements
The author would like to thank Sharvaj Kubal, prof. Yaniv Plan, and prof. Ozgur Yilmaz for many fruitful discussions. A special thanks to prof. Michael Friedlander for pointing out that a generalization of the Fenchel-Rockafellar dual would be a promising avenue of research. This paper was produced with the Latex Exporter plugin in Obsidian.
range
pages36 rangepages12 rangepages13 rangepages12 rangepages25 rangepages17 rangepages34