### v5

# Bilateral facial reduction: qualification-free subdifferential calculus and exact duality

Abstract
Qualification conditions (also termed constraint qualifications) help avoid pathological behavior at domain boundaries in convex analysis. By generalizing facial reduction from conic programming to general convex programs of the form $f(x)+g(Ax)$, we provide qualification-free generalizations of several key results: an exact Fenchel-Rockafellar dual, KKT optimality conditions, an attained infimal convolution for the conjugate of a sum, subdifferential sum and chain rules, and normal cones of intersections. All our results reduce seamlessly to their original formulations when qualification conditions hold.
The core insight is that for a sum of two convex functions, there is an affine subspace—the joint supporting subspace— that contains the feasible region, and such that qualification conditions hold when restricting the effective domain of each function to it. We offer a number of characterizations for the joint supporting subspace, including one that obtains the affine subspace via iterative, bilateral reduction between the two domains. In our proofs, which are self-contained, we develop a structured induction on faces where inductive steps are associated with normal vectors nested in supporting subspaces (a generalization of supporting hyperplanes). With this tool, we characterize the facial structure of the difference of two convex sets from the facial structures of the individual convex sets.
Keywords: facial reduction, subdifferential calculus, exact duality, qualification conditions, constraint qualifications, convex analysis
1 Introduction
In convex analysis, many results require so-called qualification conditions (sometimes called constraint qualifications), such as
$\operatorname{relint}\operatorname{dom}(f)\cap\operatorname{relint}\operatorname{dom}(g)\neq\emptyset,$ (1)
This condition can be understood as “avoiding the boundary”; it ensures that the feasible region $\operatorname{dom}(f+g)$ is not contained in the relative boundary of either $\operatorname{dom}(f)$ or $\operatorname{dom}(g)$. When Equation 1 fails to hold, the geometry of the boundary of the convex sets $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$ starts to matter, and such cases may require special treatment.
Qualification conditions are not merely theoretical artifacts; they are justified by counterexamples, and therefore reflect more or less accurately the true scope of applicability of the associated formulas. We introduce one such counterexample for the sum rule of subdifferentials. (We will later show how our theory resolves this case.) Take the function $f(x):=-\sqrt{x}$ when $x\geq 0$ and $f(x):=\infty$ otherwise. Consider the indicator function $g(x)=\iota_{(-\infty,0]}(x)$, which is 0 on $(-\infty,0]$ and $\infty$ elsewhere. Naive application of the sum rule yields
$\partial(f+g)(0)=\partial f(0)+\partial g(0)=\emptyset+\mathbb{R}_{+}=\emptyset,$
when in fact, $\partial(f+g)(0)=\partial\iota_{\{0\}}(0)=\mathbb{R}$, so the sum rule does not hold in this case. There is no contradiction in this failure because the qualification condition in Equation 1 is not satisfied in this case (otherwise, the sum rule would be guaranteed by [23, Theorem 23.8]). Indeed, the relative interiors of the domains are $(-\infty,0)$ and $(0,\infty)$, and do not intersect.
Our approach generalizes facial reduction from the conic setting to general convex programs of the form f(x)+g(Ax)f(x)+g(Ax)—let us here consider the special case $f(x)+g(x)$. We identify a single affine subspace $T_{a}\subseteq\mathbb{R}^{n}$ that simultaneously regularizes both functions and contains the feasible region. The joint supporting subspace $T_{a}$ is defined for any two convex sets (see Definition 4.1), which are usually the domains of two convex functions. We can then reduce the effective domains of each function to $T_{a}$: $f^{\prime}:=f+\iota_{T_{a}},g^{\prime}:=g+\iota_{T_{a}}$. For these reduced functions, qualification conditions are then guaranteed to hold, while preserving the original problem in the sense that $f+g\equiv f^{\prime}+g^{\prime}$. With these reduced functions we modify the sum rule to resolve our previously encountered counterexample, by using the reduced functions on the r.h.s: $\partial(f+g)(x)=\partial f^{\prime}(x)+\partial g^{\prime}(x)$. In our counter-example, the joint supporting subspace is $T_{a}=\{0\}$, and the modified sum rule yields
$\partial f^{\prime}(0)+\partial g^{\prime}(0)=\partial\iota_{\{0\}}(0)+\partial\iota_{\{0\}}(0)=\mathbb{R},$
which indeed matches $\partial(f+g)(0)$. We note that the reduced functions $f^{\prime}$, $g^{\prime}$ are trivial in this example only because the feasible region is itself trivial. In general, our reduction will preserve the functions’ behaviour inside—and to some extent outside—the feasible region.
In this manner, in Section 5 we generalize a collection of results from convex analysis, with the classical formulations trivially recovered when qualifications hold (see Proposition 5.1). Specifically, we find a modified Fenchel-Rockafellar dual for which strong duality always holds, problem-independent Karush-Kuhn-Tucker (KKT) optimality conditions, always-attained infimal convolution for the convex conjugates of sums of pairs of functions, qualification-free subdifferential sum and chain rules, and a characterization of the normal cones of the intersection of two convex sets.
We provide a number of characterizations for $T_{a}$ in Section 4. We specify the joint supporting subspace $T_{a}$ in terms of the difference of domains $\operatorname{dom}(f)-\operatorname{dom}(g)$, where the set difference is $C-D:=\{c-d\mid c\in C,d\in D\}$. Other characterizations relate $T_{a}$ to minimal faces of the domains containing the feasible region (Theorem 4.1, Corollary 4.2, and Corollary 4.3). One of our characterizations (Theorem 7.1) consists of an iterative reduction procedure which recovers $T_{a}$ and terminates in a number of steps no more than the ambient dimension. This reduction can be performed from any point in the feasible region, using so-called nested normal cones (see Definition 6.3). We show that the reduction at a point does not in fact depend on the choice of the point, which means that the reduction need only be performed once for a given optimization problem. It is known that facial reduction in the conic setting leads to a reduction of the size of the problem and improved numerical stability [7], which gives reason to believe that our generalization may have similar benefits.
A contribution of the present work is in our proof technique, found in Section 6. The central objects are nested normals (Definition 6.3) and supporting subspaces (Definition 6.2), which together describe the local lattice of faces of a convex set. We establish a structured induction via nested normals, which traverses all supporting subspaces, with induction steps associated with nested normals (see Corollary 6.1). With nested normals and our structural induction, we characterize the minimal supporting subspace containing a set (see Theorem 6.3). We also characterize the facial structure of a difference of convex sets CDC-D at 0 exclusively from the local facial structure of the individual convex sets $C$ and $D$ at $C\cap D$ (see Corollary 6.2). These results enable us to prove our characterizations of the joint supporting subspace that are presented in Section 4.
1.1 Paper outline
Section 2 surveys prior works, notably facial reduction in conic programming, limiting approaches to exact subdifferential calculus, and the lexicographic characterization of the faces of convex sets. In Section 3 we introduce some notation. In Section 4 we state analytic and constructive characterizations of the joint supporting subspace, with proofs deferred to Section 7. In Theorem 4.2 we state the main property of the joint supporting subspace: that qualification conditions hold when localizing to it. In Section 5, we state and prove generalizations of central results in convex analysis. In Section 6, we build the core theoretical framework. In Section 7, we prove the results of Section 4 with the tools of Section 6.
2 Prior works
Constraint qualifications (CQs) in optimization started with Slater’s condition [24] and much work has since been aimed at finding relaxations of these conditions [15, 1, 22]. Notably, a qualification condition using relative interiors (the weak Slater) was introduced by [23], who used the same qualification conditions for optimization problems and subdifferential calculus, which enabled our unified treatment.
There are a number of qualification-free alternatives to classical results, which come at the cost of some added complexity. One approach is to consider approximations of the core objects, with qualification-free exact results obtained in the limit of increasingly accurate approximation. This type of result includes $\varepsilon$-subdifferentials [8], and fuzzy sum rules [9, 11]. Similarly, sequential optimality conditions have been formulated [25, 10]. Sequential approaches are related to the broader theory of variational analysis.
Another kind of qualification-free result is obtained by considering the geometry of domain boundaries, with the notable example of facial reduction in conic optimization. Our work is the first, to our knowledge, to extend this geometric methodology to qualification-free subdifferential calculus.
In this vein, a qualification-free necessary and sufficient condition was formulated by [2], and extended to the nonsmooth case by [27], which achieve qualification-free characterization of minimizers by adding a term to the KKT conditions. Our qualification-free KKT condition in Corollary 5.6 is formulated for programs of the form $f(x)+g(Ax)$, instead of inequality-constrained convex programs, and uses a reduction instead of an additional term.
Our theory extends Facial Reduction (FR) from conic programs to programs of the form $f(x)+g(Ax)$. The field of facial reduction started by generalizing preprocessing ideas from linear programming to the setting of conic programming [18, 3, 5], with subsequent developments providing efficient computational methods. One computational method consists in iteratively reducing the dimension of the problem [5, 14, 26] until the minimal face is reached, an approach that we generalize in Theorem 7.1. Ideas of facial reduction can alternatively be framed as exact dual problems [4], with important applications to semidefinite programming [21, 20].
From convex geometry, a line of work that enabled the development of our theory was the lexicographical characterization of faces [17, 16]. It turned out that the lexicographical framework was not flexible enough for our needs, prompting us to develop an analogous self-contained theory, recovering the lexicographical characterization of convex faces in Theorem 6.2 in different terms. Specifically, our framework based on nested normals allows us to specify a special kind of structural induction on supporting subspaces (and thereby faces) in Corollary 6.1. This induction is powerful because each inductive step is associated with a single nested normal. Additionally, the use of nested normals allows us to specify a reduction procedure that uses nested normal cones at each step, instead of individual nested normals (see Corollary 4.4 and Theorem 7.1). This formulation provides structural insight and allows, in most cases, termination in fewer steps by dropping many dimensions at once in some steps.
Shortly after the first version of our work was made public, an independent paper by [12] was made available which specifies a reduction for convex programs of the form $f(x)+g(x)$. Their work has an algorithmic focus: they show feasibility detection and fast reductions with extended duals, by defining a notion of niceness for the domains of the convex functions. In contrast, our work is more geometric, providing analytical characterizations of the joint supporting subspace in Section 4 and qualification-free results stated with this object in Section 5, including a reduction for programs in the slightly more general form $f(x)+g(Ax)$, and subdifferential calculus sum and chain rules. Additionally, the proof techniques of the two papers seem remarkably distinct: they lift to cones, whereas we introduce nested normals in the original space.
3 Notation and Preliminaries
For two sets $A,B\subseteq\mathbb{R}^{n}$, $A+B:=\{a+b\mid a\in A,b\in B\}$ is the Minkowski sum, with $A-B:=\{a-b\mid a\in A,b\in B\}$, and for $z\in\mathbb{R}^{n}$, let $A+z:=A+\{z\}$. For a set $S\subseteq\mathbb{R}^{n}$, we denote by $\iota_{S}(x)$ the indicator function which is 0 if $x\in S$ and $\infty$ otherwise.
For a set $S\subseteq\mathbb{R}^{n}$, denote by $\operatorname{conv}(S)$ its convex hull (smallest containing convex set). For $y_{1},y_{2}\in\mathbb{R}^{n}$, let $y_{1},y_{2}]:=\operatorname{conv}\{y_{1},y_{2}\}$, and $y_{1},y_{2}[:=\operatorname{conv}\{y_{1},y_{2}\}\setminus\{y_{1},y_{2}\}$. We also denote $\operatorname{cone}(S)$ for the conical hull of a set; the smallest cone containing $S$.
Let $C\subseteq\mathbb{R}^{n}$ be convex. We denote by $\operatorname{relint}(C)$ the relative interior of $C$. A subset $F\subseteq C$ is defined to be a face of $C$ when $\forall y,z\in C$, $(\exists x\in]y,z[\cap F\implies y,z\in F)$. That is, a face $F\subseteq C$ is a subset of $C$ for which there is no line segment that “goes through” $F$ while remaining in $C$. A face $F\subseteq C$ is exposed when there is a $u\in\mathbb{R}^{n}\setminus\{0\}$ such that $F=\operatorname*{\arg\max}_{c\in C}\langle c,u\rangle$.
For any set $S\subseteq C$, we denote by $F_{C}(S)$ the face of $C$ generated by $S$; it is the smallest face of $C$ that contains $S$. (It is also the intersection of all faces containing $S$, which, it can be shown, is itself a face.)
For $x\in\mathbb{R}^{n}$ and a convex set $C\subseteq\mathbb{R}^{n}$, the normal cone of $C$ at $x$ is $N_{C}(x)=\{y\mid\forall c\in C,\langle y,c-x\rangle\leq 0\}$ if $x\in C$, and $N_{C}(x)=\emptyset$ otherwise. We denote $\mathbb{R}_{+}:=[0,\infty)$, $\mathbb{R}_{++}:=(0,\infty)$, and similarly for $\mathbb{R}_{-}$ and $\mathbb{R}_{--}$. The extended reals are $\bar{\mathbb{R}}=\mathbb{R}\cup\{-\infty,\infty\}$. Given a convex function $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ we denote by $\operatorname{dom}(f)$ the effective domain of $f$; $\operatorname{dom}(f):=\{x\in\mathbb{R}^{n}\mid f(x)\neq\infty\}$. The function $f$ is proper if it is nowhere -$\infty$ and has nonempty effective domain. The subdifferential is
$\partial f(x):=\{v\in\mathbb{R}^{n}\mid\forall y\in\mathbb{R}^{n},f(x)+\langle v,y-x\rangle\leq f(y)\},$
with $\partial f(x)=\emptyset$ when $x\notin\operatorname{dom}(f)$. The convex conjugate is $f^{*}(y):=\sup_{x\in\mathbb{R}^{n}}\langle y,x\rangle-f(x)$.
4 The joint supporting subspace
Figure 1: In each figure, the green subspace is the joint supporting subspace for the red and blue convex sets.
In this section we investigate the following object.
Definition 4.1 (The joint supporting subspace).
The joint supporting subspace of two convex sets $C,D\subseteq\mathbb{R}^{n}$ is
$T(C,D):=\operatorname{\mathrm{span}}F_{C-D}(0).$
We also define its affine counterpart: $T_{a}(C,D):=T(C,D)+C\cap D$.
We note that when $C\cap D=\emptyset$, we have $T_{a}=\emptyset$ also. This is an intentional convention which takes care of the infeasible case in Section 5. The Minkowski sum $(+C\cap D)$ in the definition of $T_{a}$ is in fact an affine translation. This follows from the next corollary.
Corollary 4.1 (Containment of the intersection).
Let $C,D\subseteq\mathbb{R}^{n}$ be convex. Then
$C\cap D-C\cap D\subseteq T(C,D).$
All proofs for results presented in this section are deferred to Section 7, because they require the machinery introduced in Section 6. Corollary 4.1 follows from Theorem 6.4; for details, the proof can be found at the end of Section 6.2.
From Corollary 4.1, we see that picking any $x\in C\cap D$, we have $T_{a}=T+x$, independently of the choice of $x$. Corollary 4.1 also implies that $C\cap D\subseteq T_{a}$, showing that $T_{a}$ is the unique translation of $T$ that contains $C\cap D$.
In this section we provide two main characterizations of the joint supporting subspace. The next result builds towards the first characterization, describing the relation of the joint supporting subspace with the individual convex sets.
Lemma 4.1 (The joint supporting subspace reveals faces).
For convex sets $C,D\subseteq\mathbb{R}^{n}$, and $T_{a}=T_{a}(C,D)$,
$T_{a}\cap C=F_{C}(C\cap D).$
Next we introduce the generated supporting subspace, which is central to our theory. This is made clear in Section 6, where we re-introduce this object in motivated fashion in Definition 6.4.
Definition 4.2 (Preliminary definition of the generated supporting subspace).
Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, define the supporting subspace of $C$ generated by $S\subseteq C$ to be
$H_{C}(S):=\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S)).$
The next theorem is our first characterization of the joint supporting subspace.
Theorem 4.1 (Characterization through generated faces).
The joint supporting subspace of $C,D\subseteq\mathbb{R}^{n}$ is
$T(C,D)=H_{C}(C\cap D)+H_{D}(C\cap D).$
Corollary 4.2 (Joint supporting subspace as hull of faces).
For convex sets $C,D\subseteq\mathbb{R}^{n}$,
$T_{a}(C,D)=\operatorname{aff}(F_{C}(C\cap D)\cup F_{D}(C\cap D)).$
Corollary 4.3 (Joint supporting subspace as difference of faces).
For convex sets $C,D\subseteq\mathbb{R}^{n}$,
$T(C,D)=\operatorname{\mathrm{span}}(F_{C}(C\cap D)-F_{D}(C\cap D)).$
So far, we have characterized the joint supporting subspace with various analytical formulas. Our next characterization instead specifies how to obtain the joint supporting subspace constructively, from an iterative process with a number of steps bounded by the ambient dimension. This result, Theorem 7.1, is best stated in terms of nested normals, a concept that we introduce in Definition 6.3, and motivate throughout Section 6. For this reason, we defer it to Section 7.3. For the moment, we state a direct corollary which does not require this object.
Corollary 4.4 (Iterative bilateral facial reduction).
Let $C,D\subseteq\mathbb{R}^{n}$ be convex sets and fix any $x\in C\cap D$. Beginning with $T_{0}=\mathbb{R}^{n}$, let
$T_{i+1}=T_{i}\cap(N_{C\cap(T_{i}+x)}(x)\cap[-N_{D\cap(T_{i}+x)}(x)])^{\perp}.$
Let $\ell$ be the smallest number such that $T_{\ell+1}=T_{\ell}$. Then:
  1. 1.
    $T_{\ell}=T(C,D)$.
  2. 2.
    $\ell\leq n$.
We defer the proof to Section 7.3. Now that the joint supporting subspace has been thoroughly identified, we move on to describing its core property. Recall that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, a common qualification condition is that for convex sets $C:=\operatorname{dom}(f),D:=\operatorname{dom}(g)$,
$\operatorname{relint}(C)\cap\operatorname{relint}(D)\neq\emptyset.$ (2)
It was shown in [23, Theorem 11.3] that this notion is equivalent to a lack of proper separation of the domains. We define the notions of separation, and proper separation below. We define these notions in affine subspaces, an intuitive generalization of separation in Euclidean spaces.
Definition 4.3 (Separation of sets).
Given an affine subspace $V\subseteq\mathbb{R}^{n}$, two convex sets $C,D\subseteq V$ are said to be separated in $V$ when there exists some $u\in(V-V)\setminus\{0\}$ such that
$\forall c\in C,d\in D,\langle u,c\rangle\leq\langle u,d\rangle.$ (3)
The concept of proper separation is slightly stronger.
Definition 4.4 (Proper separation of sets).
For an affine subspace $V\subseteq\mathbb{R}^{n}$, convex sets $C,D\subseteq\mathbb{R}^{n}$ are properly separated in $V$ when there exists some $u\in(V-V)\setminus\{0\}$ such that Equation 3 holds, and additionally such that there exists some pair $(c,d)\in C\times D$ for which the inequality in Equation 3 is strict.
Proper separation excludes the case where the two sets are merely separated by virtue of being both contained in a single hyperplane. Since separation follows from proper separation, and lack of proper separation is equivalent to intersection of relative interiors (by [23, Theorem 11.3]), the contrapositive statement is that
$\text{lack of separation}\implies\text{intersection of relative interiors}.$ (4)
In the joint supporting subspace there is always a “lack of separation” of the convex sets.
Theorem 4.2 (Qualification conditions always hold in the joint supporting subspace).
For two convex sets $C,D\subseteq\mathbb{R}^{n}$ with $T_{a}:=T_{a}(C,D)$, $C\cap T_{a}$ and $D\cap T_{a}$ are not separated in $T_{a}$.
The proof of Theorem 4.2 is the subject of Section 7.1.
Corollary 4.5 (rint qualification holds in joint supporting subspace).
For any convex sets $C,D\subseteq\mathbb{R}^{n}$ such that $C\cap D\neq\emptyset$,
$\operatorname{relint}(C\cap T_{a})\cap\operatorname{relint}(D\cap T_{a})\neq\emptyset.$
Proof of Corollary 4.5.
The result follows from Theorem 4.2 by Equation 4, with the caveat that we defined separation in affine subspaces, which are not necessarily vector spaces, while in [23, theorem 11.3] this notion is defined in $\mathbb{R}^{n}$. Yet the argument still holds because the notions of separation, proper separation, and relative interiors are invariant under translation of $T_{a}$ to $T$. For any point $x_{0}\in T_{a}$, $C\cap T_{a}$ is (properly) separated from $C\cap T_{a}$ iff $C\cap T_{a}-x_{0}$ is (properly) separated from $D\cap T_{a}-x_{0}$ in $T$. A similar equivalence holds for intersection of relative interiors. ∎
5 Main results
With Corollary 4.5 we provide qualification-free generalizations by localizing to $T_{a}$. The next result shows that when standard qualification conditions hold, the classical formulations are trivially recovered.
Proposition 5.1 (Recovery of the classical results).
If the relative interiors of two convex sets $C,D\subseteq\mathbb{R}^{n}$ intersect, then
$C\cup D\subseteq T_{a}(C,D).$
5.1 Subdifferential calculus
Corollary 5.1 (Subdifferential sum rule).
For $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ proper convex, $x\in\mathbb{R}^{n}$, and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
$\partial(f+g)(x)=\partial(f+\iota_{T_{a}})(x)+\partial(g+\iota_{T_{a}})(x).$
We note that in the infeasible case, when $f+g\equiv\infty$, $T_{a}$ is $\emptyset$, and the result holds trivially. A similar pattern occurs for many results in this section.
Proof of Corollary 5.1.
By Corollary 4.5 applied to $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$, the result follows by application of [23, Theorem 23.8] to the functions $f+\iota_{T_{a}}$ and $g+\iota_{T_{a}}$. ∎
Corollary 5.2 (Normal cone of the intersection of convex sets).
For $C,D\subseteq\mathbb{R}^{n}$ convex, and $x\in C\cap D$, for $T_{a}=T_{a}(C,D)$,
$N_{C\cap D}(x)=N_{C\cap T_{a}}(x)+N_{D\cap T_{a}}(x).$
Proof of Corollary 5.2.
With Corollary 4.5 we can use [23, Corollary 23.8.1], from which we find that
$N_{C\cap D}(x)=N_{(C\cap T_{a})\cap(D\cap T_{a})}(x)=N_{C\cap T_{a}}(x)+N_{D\cap T_{a}}(x).$
In results that follow, we abuse notation by associating a matrix $A\in\mathbb{R}^{m\times n}$ with the corresponding linear map $A:\mathbb{R}^{n}\to\mathbb{R}^{m}$. For a set $S\subseteq\mathbb{R}^{n}$, we denote by $AS$ the set image $A(S)$, and by $A^{-1}(\cdot)$ the preimage of the linear map. We make no assumption on the invertibility of the matrix $A$.
Corollary 5.3 (Subdifferential chain rule).
Let $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$ be a proper convex function, and $A\in\mathbb{R}^{m\times n}$ a matrix. Then for $T:=T(\operatorname{dom}(g),\operatorname{\mathrm{range}}(A))$,
$\partial(g\circ A)(x)=A^{T}\partial(g+\iota_{T})(Ax).$
Proof of Corollary 5.3.
Denote $T_{a}=T_{a}(\operatorname{dom}(g),\operatorname{\mathrm{range}}(A))$. Note that $\forall x\in\mathbb{R}^{n}$,
$(g\circ A)(x)=([g+\iota_{T_{a}}]\circ A)(x),$
because $T_{a}\supseteq\operatorname{dom}(g)\cap\operatorname{\mathrm{range}}(A)$. So it suffices to calculate the subdifferential of the r.h.s. of the above equation.
If $\operatorname{dom}(g)\cap\operatorname{\mathrm{range}}(A)=\emptyset,$ then the result holds trivially because $g(Ax)\equiv\infty$, and $(g+\iota_{T_{a}})(Ax)\equiv\infty$ as well. Otherwise, by Corollary 4.5, $\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})\cap\operatorname{\mathrm{range}}(A)\neq\emptyset$, which lets us apply [23, Theorem 23.9] to the composition $g+\iota_{T_{a}}]\circ A$. Additionally, note that the only face of $\operatorname{\mathrm{range}}(A)$ is itself, whence by Lemma 4.1 $\operatorname{\mathrm{range}}(A)\subseteq T_{a}$. Since $0\in\operatorname{\mathrm{range}}(A)$, this in fact implies that $T_{a}=T$, and the result follows. ∎
5.2 Infimal convolution
Recall that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the infimal convolution is
$(f\,\square g)(z):=\inf_{x,y:x+y=z}f(x)+g(y).$
In the following results, we denote by $f^{*}$ the convex conjugate of $f$.
Corollary 5.4 (Attained infimal convolution).
For two proper convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, and $T_{a}:=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
$(f+g)^{*}=(f+\iota_{T_{a}})^{*}\square(g+\iota_{T_{a}})^{*},$
where the infimum in the infimal convolution is attained.
Proof of Corollary 5.4.
Whenever $\operatorname{dom}(f)\cap\operatorname{dom}(g)=\emptyset$, we have that $T_{a}=\emptyset$, in which case the equality holds trivially, with the infimum attained at any $y\in\mathbb{R}^{n}$. Otherwise, by Corollary 4.5 the result follows from [23, Theorem 20.1]. ∎
5.3 Convex optimization
We next examine the following general convex program.
Problem 5.1 (Convex program).
Given lower semicontinuous proper convex functions $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$, and a matrix $A\in\mathbb{R}^{m\times n}$,
$\mathop{\mathrm{minimize}}_{x\in\mathbb{R}^{n}}f(x)+g(Ax).$
A standard regularity condition for this problem is that
$A(\operatorname{relint}\operatorname{dom}(f))\cap\operatorname{relint}\operatorname{dom}(g)\neq\emptyset,$
which is a weakening of Slater’s condition. A program with this property is called strongly consistent [23, Theorem 31.2].
The next result shows the sense in which 5.1 can be localized to the joint facial subspace $T_{a}$, guaranteeing strong consistency whenever the program is feasible, while retaining all the crucial information of the original program.
Theorem 5.1 (Strong consistency of the localized program).
For $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ and $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$ lower semicontinuous proper convex functions, a matrix $A\in\mathbb{R}^{m\times n}$, $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$, and any $x\in\mathbb{R}^{n}$,
$f(x)+g(Ax)=(f(x)+\iota_{A^{-1}(T_{a})}(x))+(g(x)+\iota_{T_{a}}(x)).$ (5)
And whenever $\operatorname{dom}(f+g\circ A)\neq\emptyset$,
$A\operatorname{relint}\operatorname{dom}(f+\iota_{A^{-1}(T_{a})})\cap\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})\neq\emptyset.$
Proof of Theorem 5.1.
By Corollary 4.1, $A\operatorname{dom}(f)\cap\operatorname{dom}(g)\subseteq T_{a}$, and therefore $\iota_{T_{a}}(Ax)=0$ for all $x\in\operatorname{dom}(f)\cap A^{-1}\operatorname{dom}(g)$. Additionally,
$\operatorname{dom}(f)\cap A^{-1}\operatorname{dom}(g)\subseteq A^{-1}(A\operatorname{dom}(f)\cap\operatorname{dom}(g))\subseteq A^{-1}T_{a},$
which means that $\iota_{A^{-1}T_{a}}(x)$ is 0 on $\operatorname{dom}(f+g\circ A)$. And for any $x\notin\operatorname{dom}(f+g\circ A)$, it is easily checked that both sides of Equation 5 evaluate to $\infty$, so Equation 5 holds for all $x\in\mathbb{R}^{n}$.
For the second part of the result,
$A\operatorname{relint}\operatorname{dom}(f+\iota_{A^{-1}(T_{a})})\cap\operatorname{relint}(\operatorname{dom}(g+\iota_{T_{a}}))$ (6)
$=$ $\operatorname{relint}(A\operatorname{dom}(f+\iota_{A^{-1}(T_{a})}))\cap\operatorname{relint}(\operatorname{dom}(g+\iota_{T_{a}}))$ (7)
$=$ $\operatorname{relint}(A(\operatorname{dom}(f)\cap A^{-1}(T_{a})))\cap\operatorname{relint}(\operatorname{dom}(g)\cap T_{a})$ (8)
$=$ $\operatorname{relint}(A\operatorname{dom}(f)\cap T_{a})\cap\operatorname{relint}(\operatorname{dom}(g)\cap T_{a})\neq\emptyset.$ (9)
where Equation 7 holds by [23, Theorem 6.6]; Equation 9 uses the identity that $A(S_{1}\cap A^{-1}(S_{2}))=AS_{1}\cap S_{2}$ for any sets $S_{1}\subseteq\mathbb{R}^{n}$, $S_{2}\subseteq\mathbb{R}^{m}$; and that the set in Equation 9 is non-empty is guaranteed by Corollary 4.5. ∎
We present an exact Fenchel–Rockafellar dual for 5.1: a dual problem for which strong duality always holds and the dual attained.
Corollary 5.5 (Exact Fenchel-Rockafellar dual).
For $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ and $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$ l.s.c. proper convex functions, and a matrix $A\in\mathbb{R}^{m\times n}$, let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$. Then
$\inf_{x\in\mathbb{R}^{n}}[f(x)+g(Ax)]=-\min_{y\in\mathbb{R}^{m}}[(f+\iota_{A^{-1}(T_{a})})^{*}(A^{T}y)+(g+\iota_{T_{a}})^{*}(-y)].$
Proof of Corollary 5.5.
The result follows from Theorem 5.1 with the use of [23, Corollary 31.2.1]. ∎
Definition 5.1 (Minimizer).
A minimizer of a given convex program is any point $x\in\mathbb{R}^{n}$ that attains the infimum of the program, provided that the infimum is finite.
Next, we provide a characterization of the minimizers of 5.1, i.e., KKT optimality conditions. Our KKT condition holds without any further assumptions on the program; we do not require strong consistency.
Corollary 5.6 (KKT conditions).
Consider 5.1, and let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$. A point $x\in\mathbb{R}^{n}$ is a minimizer of 5.1 if, and only if, there exists some $y\in\mathbb{R}^{m}$ such that
  1. 1.
    $Ax\in\partial(g+\iota_{T_{a}})^{*}(y).$
  2. 2.
    $A^{T}y\in\partial(f+\iota_{A^{-1}(T_{a})})(x)$.
Proof of Corollary 5.6.
From Theorem 5.1, the result follows from application of [23, Corollary 31.3.1] to the functions $f+\iota_{A^{-1}(T_{a})}$ and $g+\iota_{T_{a}}$. ∎
6 A framework for local face lattices
Inspired by the lexicographic characterization of convex faces of [16], we develop a framework describing the lattice of convex faces at a point $x\in C$, which we will then use to show the results of Section 4. We begin with a characterization of faces of convex sets. In the next few results, we denote $X$ for a real vector space.
Proposition 6.1 (Characterization of convex faces by supporting subspaces).
Let $C\subseteq X$ be a convex set. A subset $F\subseteq C$ is a face of $C$ if, and only if the two following conditions hold.
  1. 1.
    $\operatorname{aff}(F)\cap C=F$,
  2. 2.
    $C\setminus F$ is convex.
Proof of Proposition 6.1.
Let $F$ be a face. We argue by contradiction that $\operatorname{aff}(F)\cap[C\setminus F]=\emptyset$; let us assume that there is some $z\in\operatorname{aff}(F)\cap[C\setminus F]$.
If $F$ is a singleton, then $\operatorname{aff}(F)=F$, and $\operatorname{aff}(F)\cap[C\setminus F]=\emptyset$. Otherwise, we may assume that $0\in F\setminus\{z\}$ without loss of generality, because the faces of a convex set are invariant under translation. With this assumption, $\operatorname{aff}(F)=\operatorname{\mathrm{span}}(F)=\operatorname{cone}(F-F)$. Then $z\in\operatorname{cone}(F-F)$, so there are $y_{1},y_{2}\in F$ such that $z=\alpha(y_{1}-y_{2})$ for some $\alpha\geq 0$. We can rule out the possibility that $z\in[y_{1},y_{2}]$, because by convexity of $F$ it would follow that $z\in F$. The remaining possibility is that $y_{1}\in[y_{2},z[$. Further, note that $y_{1}\neq y_{2}$, because otherwise $z=0$, which disagrees with our assumption that $z\in F\setminus\{0\}$. Therefore, $y_{1}\in]y_{2},z[$. The line segment $y_{2},z]$ has a point $y_{1}\in]y_{2},z[\cap F$. Since $y_{2},z\in C$, the fact that $F$ is a face implies that $z$ must then be in $F$, in contradiction with our definition of $z$. We have therefore shown the claim.
We next argue that $C\setminus F$ is convex. $\forall y,z\in C\setminus F$, any $x\in]y,z[$ is in $C$ by convexity of $C$. Also, $x$ is not in $F$ because $F$ is a face. So any $x\in]y,z[$ is also in $C\setminus F$, which means that $C\setminus F$ is convex.
We now show the converse: let $F\subseteq C$ be such that $C\setminus F$ is convex and $\operatorname{aff}(F)\cap C=F$, we show that $F$ is a face of $C$.
Take any $y,z\in C$, $x\in F\cap]y,z[$. If $y\in\operatorname{aff}(F)$ then $z\in\operatorname{aff}(F)$ as well, because $z=\alpha(x-y)+y$ for some $\alpha>0$. So either $z,y$ are both in $\operatorname{aff}(F)$, or neither are. But if neither are, then $z,y\in C\setminus\operatorname{aff}(F)$, and by convexity of $C\setminus\operatorname{aff}(F)$, $x\in C\setminus\operatorname{aff}(F)$ also, which is a contradiction with $x\in F$. Therefore, $z,y$ are both in $F$, which shows that $F$ is a face. ∎
Proposition 6.1 motivates the next mathematical object.
Definition 6.1 (Supporting subspace).
Let $C\subseteq X$ be convex. A supporting subspace $L\subseteq X$ of $C$ is an affine subspace such that $C\setminus L$ is convex.
Though supporting subspaces are affine subspaces, we omit the adjective “affine” for brevity. This notion appears in [6, Exercise 5.4], and is related to order ideals, a concept defined for convex cones. Supporting subspaces are of central importance to this paper; it allows us to shift our focus away from the faces themselves to the comparatively simpler affine subspaces which select them through intersections with the convex set. The next result describes the relationship between supporting subspaces and faces.
Proposition 6.2 (Relation of supporting subspaces to faces).
Let $C\subseteq X$ be convex.
  1. 1.
    For any affine subspace $L\subseteq X$,
[…]
  2. 2.
    For any subset $F\subseteq C$,
[…]
Proof of Proposition 6.2.
1) ($\implies$) We show that if $L$ is a supporting subspace, then $F:=C\cap L$ satisfies both conditions of Proposition 6.1:
  1. 1.
    That $\operatorname{aff}(F)\cap C=F$: note that $\operatorname{aff}(F)\subseteq L$ because $F\subseteq L$, and so $\operatorname{aff}(F)\cap C\subseteq L\cap C=:F$. On the other hand, $L\cap C=(L\cap C)\cap C\subseteq\operatorname{aff}(L\cap C)\cap C$, and so $F\subseteq\operatorname{aff}(F)\cap C$.
  2. 2.
    That $C\setminus F$ is convex: note that
    $C\setminus F=C\setminus[C\cap L]=C\setminus L,$
    which is convex by definition of $L$ as a supporting subspace.
[…]
2) ($\implies$) If $F$ is a face, then we argue that $\operatorname{aff}(F)$ is a suitable supporting subspace. Indeed, by Proposition 6.1 we know that $F=C\cap\operatorname{aff}(F)$ and that $C\setminus F$ is convex. Note that $C\setminus F=C\setminus(C\cap\operatorname{aff}(F))=C\setminus\operatorname{aff}(F)$, and therefore $C\setminus\operatorname{aff}(F)$ is convex, while $F=C\setminus\operatorname{aff}(F)$, which shows that $\operatorname{aff}(F)$ is a supporting subspace that selects $F$ by intersection.
[…]
In this work, we focus on “local” supporting subspaces; the subset of supporting subspaces that contain a point $x\in C$. These subsets have some additional nice algebraic properties. We find it useful to generalize this idea of “locality” to supporting subspaces that are localized at a set $S\subseteq C$ instead of only at points, i.e., we consider all supporting subspaces of $C$ that also contain $S$. To emphasize this connection, we say that a supporting subspace is “at SS” to describe its containment of $S$. The reader should keep in mind the important special case of $S$ being a singleton $S=\{x\}$, for which the results are localized at a point.
Any supporting subspace $L$ at $S$ has its affine offset determined by $S$. Recall that LLL-L is the tangent space of $L$, i.e., the translation of $L$ that becomes a (non-affine) subspace. With this, it becomes clear that L=(LL)+SL=(L-L)+S. We can always deduce the correct affine offset from $S$ alone. Therefore, we find it convenient to describe the set of supporting subspaces that contain $S$ through their tangent spaces. We henceforth call the tangent spaces of supporting subspaces “supporting subspaces” as well, and whether we mean the affine subspace or its tangent space will be clear from context.
Definition 6.2 (Supporting subspace at S).
Let $C\subseteq X$ be a convex set and $S\subseteq X$ a non-empty subset. We call a subspace $U\subseteq X$ a supporting subspace of $C$ at $S$ if:
  1. 1.
    $S-S\subseteq U$
  2. 2.
    $U+S$ is an (affine) supporting subspace of $C$ as defined in Definition 6.1.
We define $\mathcal{F}(C;S)$ to be the set of all subspaces $U$ satisfying both conditions above for fixed $C$ and $S$.
When $S$ is the singleton $\{x\}$, the requirement that $S-S\subseteq U$ is trivially satisfied. When $S$ is not a singleton, this assumption guarantees that $\forall s_{1},s_{2}\in S,U+s_{1}=U+s_{2}$. This implies that $\forall s\in S,U+s=U+S$, and so the Minkowski sum $U+S$ is a simple affine offset of $U$, for which $S\subseteq U+S$. In other words, the supporting subspace $U+S$ contains $S$. We introduce the notion of a nested normal cone. Let $C\subseteq\mathbb{R}^{n}$ be convex and $U\subseteq\mathbb{R}^{n}$ be a subspace, then $\forall x\in\mathbb{R}^{n}$,
$N_{C}^{U}(x):=N_{C\cap(x+U)}(x)\cap U.$
We would also like for our normal cones to located at a set $S\subseteq C$, similarly to supporting subspaces.
$N_{C}(S):=\{v\in\mathbb{R}^{n}\mid\forall s\in S,c\in C,\langle c-s,v\rangle\leq 0\}.$
The hyperplanes corresponding to the vectors in this normal cone are supporting hyperplanes of $C$ that contain $S$. Note also that $N_{C}(S)=\bigcap_{s\in S}N_{C}(s).$ The idea of normal cones that are located at sets appears in the earlier work [13], where normal cones are defined for faces of polytopes, instead of at points. Combining the two generalizations, we get the following.
Definition 6.3 (Nested normals).
Let $C\subseteq\mathbb{R}^{n}$ be a convex set, and $S$ a non-empty subset of $C$. Then for a subspace $U\subseteq\mathbb{R}^{n}$ such that $S-S\subseteq U$, and with $C^{\prime}:=C\cap(U+S)$,
$N_{C}^{U}(S):=\{v\in U\mid\forall c\in C^{\prime},s\in S,\langle c-s,v\rangle\leq 0\}$
is a nested normal cone of $C$ at $S$, and we label any $v\in N_{C}^{U}(S)$ a nested normal of $C$ at $S$ in $U$.
The normal cone at a set has an elegant connection to the notion of relative interior of that set: the normal cone at points in the relative interior match the normal cone of the set.
Lemma 6.1 (Normal cones in the relative interior are representative).
Let $C\subseteq\mathbb{R}^{n}$ be convex and $S\subseteq C$ a non-empty convex set. Let $U\subseteq\mathbb{R}^{n}$ be a subspace such that $S-S\subseteq U$, and $x\in\operatorname{relint}(S)$. Then
$N_{C}^{U}(S)=N_{C}^{U}(x).$
Proof of Lemma 6.1.
That $N_{C}^{U}(S)\subseteq N_{C}^{U}(x)$ follows from simple inclusion $x\in S$.
For the containment ($\supseteq$), WLOG let $U=\mathbb{R}^{n}$. We can make this assumption because the normal cones are invariant under the translation $U+S\to U$, and so we have an equivalent normal cone inside the subspace $U$. Since $U$ is a finite dimensional vector space, it is isomorphic to $\mathbb{R}^{n}$ for some $n\in\mathbb{N}$.
Let $v\in N_{C}(x)$. Then $\forall s\in S,\langle v,s\rangle\leq\langle v,x\rangle$ because $S\subseteq C$. We show that this inequality is in fact always an equality. Arguing by contradiction, assume that there is an $s\in S$ such that $\langle v,s\rangle<\langle v,x\rangle$. Then since $x\in\operatorname{relint}(S)$, there is a $\varepsilon>0$ such that $(\varepsilon B+x)\cap\operatorname{aff}(S)\subseteq S$ for $B$ the unit ball in $\mathbb{R}^{n}$. There is a $\lambda>0$ such that $s^{\prime}:=x+\lambda(s-x)\in\varepsilon B+x$. Then note that $s^{\prime}\in(\varepsilon B+x)\cap\operatorname{aff}(S)\subseteq S$, and
$\langle v,s^{\prime}\rangle=\langle v,x\rangle-\lambda\langle v,s-x\rangle.$
The second term is strictly positive, therefore $\langle v,s^{\prime}\rangle>\langle v,x\rangle$. But this is a contradiction with $v\in N_{C}(x)$. Therefore, we find that $\forall s\in S,$ $\langle v,s\rangle=\langle v,x\rangle$. Then $\forall c\in C,s\in S$ $\langle c-s,v\rangle=\langle c-x,v\rangle\leq 0$, and therefore $v\in N_{C}(S)$. ∎
Next, we show that the nested normals at a set $S$ are orthogonal with SSS-S.
Lemma 6.2 (Normal cone everywhere is the orthogonal subspace).
Let $C\subseteq\mathbb{R}^{n}$ be a convex set, and $U\subseteq\mathbb{R}^{n}$ a subspace. Then for any non-empty $S\subseteq C$ such that $S-S\subseteq U$,
$N_{C}^{U}(S)\subseteq(S-S)^{\perp}\cap U.$
Further, when $S=C\cap(U+S)$, the sets are equal.
Proof of Lemma 6.2.
Let $y\in N_{C}^{U}(S)$. By Definition 6.3, $\forall z,x\in S,$ since $x\in C$ and $z\in S$, $\langle x-z,y\rangle\leq 0,$ and since $x\in S$ and $z\in C$, $\langle x-z,y\rangle\geq 0$. Then $y\perp S-S$, which shows the inclusion.
Next we show the containment ($\supseteq$), in the case that $S=C\cap(U+S)$. Let $y\in(S-S)^{\perp}\cap U$. For any $x\in S,z\in C\cap(U+S)$ we need to show that $\langle z-x,y\rangle\leq 0$. Note that $z\in S$, so $x-z\in S-S$, and by definition of $y$, $\langle z-x,y\rangle=0$, which proves the statement. ∎
The core property of nested normals is that they describe the lattice structure of supporting subspaces, which is closely related to the lattice structure of the faces. We make this idea precise in the next two results.
Theorem 6.1 (Nested normals induce supporting subspaces).
Let $C\subseteq\mathbb{R}^{n}$ be convex, $S\subseteq C$ non-empty, $U\subseteq\mathbb{R}^{n}$ a subspace, and $v\in N_{C}^{U}(S)$. Let $V\in\mathcal{F}(C;S)$ be such that $V\subseteq U$. Then
$V\cap v^{\perp}\in\mathcal{F}(C;S).$
Proof of Theorem 6.1.
First, note that if $v=0$, then $V\cap v^{\perp}=V$, and there result holds. We now consider the case where $v\neq 0$.
By Lemma 6.2, $v\perp S-S$, and therefore $S-S\subseteq V\cap v^{\perp}$.
We make two observations. First, since $v\in N_{C}^{U}(S)$ and $V\subseteq U$, $\forall y\in C\cap(V+S)$, $\forall s\in S,$ we have that $\langle y,v\rangle\leq\langle s,v\rangle$. Second, this inequality is tight precisely for those values of $y$ that are also contained in $v^{\perp}+S$, i.e.,
$\{y\in C\cap(V+S)\mid\forall s\in S,\langle y,v\rangle=\langle s,v\rangle\}$
$=$ $((V+S)\cap C)\cap(v^{\perp}+S)=(V\cap v^{\perp}+S)\cap C.$
These two facts imply that
$(V\cap v^{\perp}+S)\cap C=\operatorname*{\arg\max}_{y\in C\cap(V+S)}\langle y,v\rangle.$
The r.h.s. is an exposed face of the face $C\cap(V+S)$, so it is itself a face, which by Proposition 6.2 implies that $V\cap v^{\perp}\in\mathcal{F}(C;S)$. ∎
We next show that all supporting subspaces are “reached” by nested normals in supporting subspaces.
Theorem 6.2 (Completeness of the composition of nested normals).
Let $C\subseteq\mathbb{R}^{n}$ be convex, $S\subseteq C$ non-empty, and $V,U\in\mathcal{F}(C;S)$ such that $U\subseteq V$. Then there is a sequence $V_{0},\ldots,V_{k}\in\mathcal{F}(C;S)$ with $k=\dim(V)-\dim(U)$ such that $V_{0}=V$, $V_{k}=U$, and for each $i\in\{1,\ldots,k\}$ there is a $v_{i}\in N_{C}^{V_{i-1}}(S)\setminus\{0\}$ such that $V_{i}=V_{i-1}\cap v_{i}^{\perp}$.
This result is similar to the lexicographic characterization of the faces of convex sets in [16, Proposition 6], which states that every face can be obtained by a sequence of faces where each face is an exposed face of the preceding face. Let us state a technical lemma that we will need for the proof.
Lemma 6.3 (Affine slices preserve supporting subspaces).
Let $C\subseteq\mathbb{R}^{n}$ be convex and $S\subseteq C$ non-empty. Let $U\in\mathcal{F}(C;S)$, and $V$ any subspace such that $S-S\subseteq V$. Then $U\cap V\in\mathcal{F}(C\cap(V+S);S)$.
Proof of Lemma 6.3.
Since $C\setminus(U+S)$ is convex, $(V+S)\cap(C\setminus(U+S))$ is also convex. We can write this set as $(V+S)\cap C]\setminus[(V+S)\cap(U+S)]$, which is the same as $((V+S)\cap C)\setminus((V\cap U)+S)$, because $S-S\in V\cap U$. Since this set is convex, and $S-S\subseteq V\cap U$, we find that $V\cap U\in\mathcal{F}((V+S)\cap C;S)$. ∎
Proof of Theorem 6.2.
It suffices to show the following statement: given any $U,V\in\mathcal{F}(C;S)$ such that $V\supsetneq U$, there exists a $v\in N_{C}^{V}(S)\setminus\{0\}$ such that $v\perp U$. If this statement is shown to hold, applying it recursively generates the sequence of supporting subspaces $V_{1},\ldots,V_{k}$ in a finite number of steps. Note that the dimension of the supporting subspaces decreases by one at each iteration while always containing $U$, which implies that $U\subseteq V_{k}$ with $\dim(U)=\dim(V_{k})$, and therefore $U=V_{k}.$
We now show the above statement. Let $C^{\prime}=C\cap(V+S)$, $\bar{C}=\operatorname{\Pi}_{U^{\perp}}C^{\prime}$, and $\bar{U}=\operatorname{\Pi}_{U^{\perp}}(U+S)$. Note that $\bar{C}$ is a convex set and $\bar{U}$ a singleton $\{x\}\subseteq\bar{C}$. We argue that $\bar{U}$ is a face of $\bar{C}$. Arguing by contradiction, we hypothesize the $\bar{U}$ is not a face of $\bar{C}$, in which case there are $\bar{y},\bar{z}\in\bar{C}\setminus\bar{U}$ such that $x=\lambda\bar{y}+(1-\lambda)\bar{z}$ for $\lambda\in(0,1)$. This implies that there are elements $y,z\in C^{\prime}$ such that $\bar{y}=\operatorname{\Pi}_{U^{\perp}}y$ and $\bar{z}=\operatorname{\Pi}_{U^{\perp}}z$, with $w=\lambda y+(1-\lambda)z$. Note that the line segment $[y,z]$ has endpoints $y,z\in C^{\prime}\setminus(U+S)$ with a point $w\in]y,z[$ that is also contained in $C^{\prime}\cap(U+S)$. This implies that $C^{\prime}\cap(U+S)$ is not a face of $C^{\prime}$, which, by Proposition 6.2, means that $U\notin\mathcal{F}(C^{\prime};S)$. This is a contradiction because in fact, $U\in\mathcal{F}(C^{\prime};S)$, as we now show. We know that $U\in\mathcal{F}(C;S)$, so by Lemma 6.3, $U=U\cap V\in\mathcal{F}(C\cap(V+S);S)$, i.e., $U\in\mathcal{F}(C^{\prime};S)$. Therefore, we indeed have a contradiction, from which we find that $\bar{U}$ is a face of $\bar{C}$.
Note that $\operatorname{\Pi}_{U^{\perp}}(V+S)\subseteq V+S$. This becomes apparent from the following equality: $V+S=(V+S)+U=\operatorname{\Pi}_{U^{\perp}}(V+S)+U,$ where the first equality holds because $U\subseteq V$ implies that $V=V+U$.
Then since $U+S,C\cap(V+S)\subseteq V+S$, we find that $\bar{U},\bar{C}\subseteq(V+S)\cap U^{\perp}$. Since $\bar{U}$ is a face of $\bar{C}$, there exists a supporting hyperplane $H$ of $\bar{C}$, where the hyperplane is relative to the affine subspace $(V+S)\cap U^{\perp}$, such that $\bar{U}\subseteq H$. Indeed, if there is no such hyperplane $H$, $\bar{U}$ would then be a singleton in the interior of $\bar{C}$ (where the interior is relative to $(V+S)\cap U^{\perp}$, i.e., with the induced topology in $(V+S)\cap U^{\perp}$).
Take $\bar{v}\in V\cap U^{\perp}$ to be the unit normal to $H$ that is also in $N_{\bar{C}}^{V\cap U^{\perp}}(x)$. Then note that $\bar{v}\in N_{C}^{V}(S)\cap U^{\perp}$, which is the vector we had to find. ∎
Together, Theorem 6.1 and Theorem 6.2 enables a powerful tool that is central to our proofs: we can perform structural induction on the tree of supporting subspaces, using only nested normals. Indeed, the set of supporting subspaces forms a tree, where the nodes are the supporting subspaces, the root is $\mathbb{R}^{n}$ (note that it is always a supporting subspace, and that it contains any other supporting subspace), and every edge of the tree is associated with a nested normal. Crucially, Theorem 6.2 shows that starting from the root, and generating the tree by using the nested normals, we attain all supporting subspaces. Induction on this tree then takes the following form.
Corollary 6.1 (Induction by nested normal).
Let $C$ be a convex set, $S\subseteq C$ a non-empty subset, and a statement $A:\mathcal{F}(C;S)\to\{\text{true},\text{false}\}$. If the following statements hold:
  1. 1.
    $A(\mathbb{R}^{n}),$
  2. 2.
    $\forall U\in\mathcal{F}(C;S)$, $v\in N^{U}_{C}(S)$, $A(U)\implies A(U\cap v^{\perp})$, then it follows that
    $\forall V\in\mathcal{F}(C;S),\,A(V).$
Proof of Corollary 6.1.
Consider any $U\in\mathcal{F}(C;S)$, and let the statement $A$ satisfy both stated conditions. Then by Theorem 6.2, we have supporting subspaces $V_{0},\ldots,V_{k}\in\mathcal{F}(C;S)$ and nested normals $v_{1},\ldots,v_{k}$ as in Theorem 6.2. $A(V_{0})=\mathrm{true}$ because $V_{0}=\mathbb{R}^{n}$. Also, $\forall i\in[k]$, $A(V_{i-1})\implies A(V_{i})$ because of the existence of the nested normal $v_{i}\in N_{C}^{V_{i-1}}(S)\setminus\{0\}$ with the property that $V_{i}=V_{i-1}\cap v_{i}^{\perp}$. Then $A(U)$ holds by classical induction, because $U=V_{k}$. ∎
6.1 Minimal supporting subspace
Recall that the smallest face of a convex set $C$ containing some $x\in C$ is $F_{C}(x)$, the face of $C$ generated by $x$. It is the only face with $x$ in its relative interior. In this section, we develop an analogous notion for supporting subspaces, and characterize this supporting subspace by using nested normals.
Definition 6.4 (The generated supporting subspace).
Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, we define the generated supporting subspace $H_{C}(S)$ of $C$, generated by $S$, to be the supporting subspace that is contained in every other supporting subspace in $\mathcal{F}(C;S)$.
We show existence and uniqueness for the above definition, starting with a lemma.
Lemma 6.4 (Supporting subspaces are closed under intersections).
For $C\subseteq\mathbb{R}^{n}$ convex and $S\subseteq C$ non-empty, let $U,V\in\mathcal{F}(C;S)$. Then $U\cap V\in\mathcal{F}(C;S)$.
Proof of Lemma 6.4.
The subspace $U\cap V$ is such that
$(U\cap V+S)\cap C=((U+S)\cap C)\cap((V+S)\cap C).$
The r.h.s. is the intersection of two convex faces of $C$. Their intersection is then also a face (immediate from the definition of faces). By Proposition 6.1, $C\setminus[(U\cap V+S)\cap C]$ is convex. The convexity of this set, which we can write as $C\setminus((U\cap V)+S)$, implies that $U\cap V\in\mathcal{F}(C;S)$. ∎
Proposition 6.3.
The generated supporting subspace of Definition 6.4 exists and is unique.
Proof of Proposition 6.3.
For existence, starting at any supporting subspace $U\in\mathcal{F}(C;S)$, either $U$ contains no other supporting subspace in $\mathcal{F}(C;S)$, or we have a $U^{\prime}\in\mathcal{F}(C;S)$ such that $U^{\prime}\subsetneq U$. Repeating this argument, we eventually find a supporting subspace that contains no other, because the dimension is reduced at every step, and the ambient dimension is finite.
For uniqueness, let $V_{1},V_{2}\in\mathcal{F}(C;S)$ be such that they properly contain no other supporting subspace in $\mathcal{F}(C;S)$. Then by Lemma 6.4, $V_{1}\cap V_{2}\in\mathcal{F}(C;S)$ is properly contained in both $V_{1}$ and $V_{2}$, which is a contradiction. ∎
As might be expected, the generated supporting subspace is closely related to the generated face.
Lemma 6.5 (Relation between the generated face and the generated supporting subspace).
With $C\subseteq\mathbb{R}^{n}$ convex and $S\subseteq C$ non-empty,
  1. 1.
    $F_{C}(S)=C\cap(H_{C}(S)+S).$
  2. 2.
    $H_{C}(S)=\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S)).$
Proof of Lemma 6.5.
By Proposition 6.1, $H_{C}(S)$ being the smallest supporting subspace implies that $C\cap(H_{C}(S)+S)$ is the smallest face containing $S$, because any properly contained, smaller face would have its own supporting subspace (by Proposition 6.2), which would have to be properly contained in $H_{C}(S)$.
For the second statement, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is a supporting subspace because $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))+S=\operatorname{aff}(F_{C}(S))$, which is a supporting subspace, as implied by Proposition 6.1. Further, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is contained in any other supporting subspace in $\mathcal{F}(C;S)$, because given $U\in\mathcal{F}(C;S)$ such that $U\subsetneq\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$, we would then have $C\cap(U+S)$ as a face of $C$ containing $S$ that is properly contained in $F_{C}(S)$, in contradiction with the definition of $F_{C}(S)$. ∎
We can now restate Definition 4.1 with specialized terms.
Definition 6.5 (Reformulated definition of the joint supporting subspace).
The joint supporting subspace of two convex sets $C,D\subseteq\mathbb{R}^{n}$, as defined in Definition 4.1, is
$T(C,D):=H_{C-D}(0).$
The next result characterizes generated supporting subspaces in terms of nested normals.
Theorem 6.3 (The generated supporting subspace is the only supporting subspace with no nested normals).
Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, the generated supporting subspace $H_{C}(S)$ is the only element in $\mathcal{F}(C;S)$ such that $N_{C}^{H_{C}(S)}(S)=\{0\}$.
Proof of Theorem 6.3.
The generated supporting subspace does not contain any non-trivial nested normals, because any non-trivial nested normal $v\in N_{C}^{H(C)}(S)\setminus\{0\}$ induces the supporting subspace $H_{C}(S)\cap v^{\perp}\subsetneq H_{C}(S)$ by Theorem 6.1.
For any supporting subspace $V\in\mathcal{F}(C;S)$ such that $N_{C}^{V}(S)=\{0\}$, we argue that $V=H_{C}(S)$. Note that $H_{C}(S)\subseteq V$ by Definition 6.4. The inclusion cannot be proper, because otherwise, by Theorem 6.2, there would be a nontrivial nested normal $v\in N_{C}^{V}(S)\setminus\{0\}$. ∎
6.2 Nested normals of a difference of sets
We characterize the supporting subspaces of a difference of sets CDC-D from the supporting subspaces and nested normals of the individual sets $C$ and $D$ at $C\cap D$. One might hope for the simple statement that $\mathcal{F}(C;S)\cap\mathcal{F}(D;S)=\mathcal{F}(C-D;0)$ for $S\subseteq C\cap D$. Unfortunately, though $(\supseteq)$ turns out to hold, $(\subseteq)$ does not in general.
Example 6.1 (Mismatched nested normals).
Let $h:\mathbb{R}\to\mathbb{R},h(x):=x^{2}$, $C:=\operatorname{epi}\min(\iota_{\mathbb{R}_{-}},h)$ and $D:=\mathrm{hypo}\max(-\iota_{\mathbb{R}_{-}},-h)$, where $\mathrm{hypo}$ denotes the hypograph of the concave function. Then $\{0\}$ is a face of both $C$ and $D$, yet $C-D=\mathbb{R}\times\mathbb{R}_{+}$, and therefore $\{0\}\notin\mathcal{F}(C-D;0)$.
There is therefore a need for some additional condition on supporting subspaces shared by $C$ and $D$ for them to also be supporting subspaces of CDC-D. The next result allows us to derive such a condition by using nested normals.
Theorem 6.4 (Decomposition of supporting subspaces of differences).
Let $C,D\subseteq\mathbb{R}^{n}$ be convex with $S\subseteq C\cap D$ non-empty, and $U\in\mathcal{F}(C-D;0)$. Then
  1. 1.
    $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S).$
  2. 2.
    $N^{U}_{C-D}(0)=N^{U}_{C}(S)\cap(-N^{U}_{D}(S)).$
  3. 3.
    $C\cap(U+S)-D\cap(U+S)=(C-D)\cap U.$
The next corollary shows that the missing condition is the existence of a synchronized “path” of nested normals for $C$ and $D$.
Corollary 6.2 (Characterization of the faces of differences of convex sets).
Let $C,D\subseteq\mathbb{R}^{n}$ be convex and $S\subseteq C\cap D$ non-empty. The elements of $\mathcal{F}(C-D;0)$ are precisely the elements $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$ for which there exists a sequence $V_{0},\ldots,V_{k}\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$ such that $V_{0}=\mathbb{R}^{n}$, $V_{k}=U$, and such that for each $i\in\{1,\ldots,k\}$ there is a $v_{i}\in N_{C}^{V_{i-1}}(S)\cap-N_{D}^{V_{i-1}}(S)\setminus\{0\}$ such that $V_{i}=V_{i-1}\cap v_{i}^{\perp}$.
Proof of Corollary 6.2.
That any $U$ with such a sequence is in $\mathcal{F}(C-D;0)$ follows by first noticing that each $v_{i}$ in the sequence belongs to $N^{U}_{C-D}(0)$ by Theorem 6.4. By applying Theorem 6.1 sequentially, it follows that $U\in\mathcal{F}(C-D;0)$.
That any element $U\in\mathcal{F}(C-D;0)$ has such a sequence follows from Theorem 6.2, which shows that there is a sequence of nested normals of CDC-D at 0 that, starting from $\mathbb{R}^{n}$, attains $U$. By Theorem 6.4, this sequence corresponds to a sequence $v_{i},\ldots,v_{k}$ as described in the statement. ∎
We can now better understand Example 6.1. The supporting subspace $\{0\}$ of $C$ and $D$ is obtained via the sequence of nested normals $\{(0,-1),(1,0)\}$ for $C$ and $\{(0,1),(1,0)\}$ for $D$. Only the first nested normals are opposite to each other, and therefore there is no synchronized path of nested normals to $\{0\}$. If we modify the example by taking $D:=\mathrm{hypo}\max(-\iota_{\mathbb{R}_{+}},-h)$, then $\{0\}$ is attained by $\{(0,1),(-1,0)\}$, and the sequence is synchronized with that of $C$. In this modified example, $\{0\}$ is indeed a supporting subspace of CDC-D because CDC-D is a convex set such that $C-D\subseteq\mathbb{R}\times\mathbb{R}^{+}$ and $C-D\cap(\mathbb{R}\times\{0\})=\mathbb{R}_{-}\times\{0\}$, which shows that $\{0\}\in\mathcal{F}(C-D;0)$.
To show Theorem 6.4, we require the following lemma.
Lemma 6.6 (Normal cone of the difference of convex sets).
Given convex sets $C,D\subseteq\mathbb{R}^{n}$ with $S\subseteq C\cap D$ non-empty,
$N_{C-D}(0)=N_{C}(S)\cap(-N_{D}(S)).$
Proof of Lemma 6.6.
We start with the inclusion $(\subseteq)$. Any $v\in N_{C-D}(0)$ has the property that $\forall c\in C,d\in D$, $\langle v,(c-d)-0\rangle\leq 0$. Then $\forall c\in C,s\in S$, $\langle v,c-s\rangle\leq 0$ because $s\in D$, and therefore $v\in N_{C}(S)$. By a similar argument, $v\in-N_{D}(S)$, which shows the inclusion.
For the containment ($\supseteq$), fix some $s\in S$. let $v\in N_{C}(S)\cap(-N_{D}(S))$. Then $\forall c\in C,d\in D$
$\langle v,c-d\rangle=\langle v,c-s\rangle+\langle v,s-d\rangle\leq 0,$
and so $v\in N_{C-D}(0)$. ∎
Proof of Theorem 6.4.
We argue by nested normal induction (from Corollary 6.1) over the supporting subspaces of CDC-D. Fixing some $S\subseteq C\cap D$, let $A(U)$ be the statement that all three points in the statements hold for supporting subspace $U\in\mathcal{F}(C-D;0)$.
The base case, where $U=\mathbb{R}^{n}$, holds because $\mathbb{R}^{n}$ is always a supporting subspace. Point 2) is given by Lemma 6.6, and point 3) holds trivially.
For the inductive step, let $V\in\mathcal{F}(C-D;0)$ be such that $A(V)$ is true. Let $v\in N_{C-D}^{V}(0)\setminus\{0\}$, and $U=V\cap v^{\perp}$. In the rest of the proof, we show that $A(U)$ is true. Note that by point 2) of $A(V)$, $v\in N_{C}^{V}(S)\cap(-N_{D}^{V}(S))$, which by Theorem 6.1 implies that $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$, and this shows point 1) of $A(U)$.
We next show that for $C^{\prime}=C\cap(V+S)$, $D^{\prime}=D\cap(V+S)$,
$(C^{\prime}-D^{\prime})\cap U=(C^{\prime}\cap(U+S))-(D^{\prime}\cap(U+S)).$ (10)
We start with the inclusion ($\subseteq$): Take $z\in C^{\prime}$, $y\in D^{\prime}$ with $z-y\in U$. Then since $U\subseteq v^{\perp}$, $\langle z-y,v\rangle=0$, and therefore $\forall s\in S$, $\langle z-s,v\rangle-\langle y-s,v\rangle=0$. That $\langle z-s,v\rangle\leq 0$ follows from $v\in N_{C}^{V}(S)$ since $z\in C^{\prime}\subseteq C$ and $z-s\in V$, and similarly, $\langle y-s,v\rangle\geq 0$ because $v\in N_{D}^{V}(S)$. But then the only way that $\langle z-s,v\rangle-\langle y-s,v\rangle=0$ is if $\langle z-s,v\rangle=0$ and $\langle y-s,v\rangle=0$ individually. It follows that $z,y\in U+S$, and so the inclusion holds. The containment ($\supseteq$) is trivial, so we have shown Equation 10.
By point 3) of $A(V)$, $C^{\prime}-D^{\prime}=(C-D)\cap V$. Also, since $U\subseteq V$, $C^{\prime}\cap(U+S)$ = $C\cap(U+S)$, and similarly for $D$. Applying these identities on both sides of Equation 10, we find that
$(C-D)\cap U=C\cap(U+S)-D\cap(U+S).$
This shows that point 3) of $A(U)$ is true.
To show point 2) of $A(U)$, we compute
$N_{C}^{U}(S)\cap(-N_{D}^{U}(S))$ $=[N_{C\cap(U+S)}(S)\cap(-N_{D\cap(U+S)}(S))]\cap U$ (11)
$=N_{C\cap(U+S)-D\cap(U+S)}(0)\cap U$ (12)
$=N_{(C-D)\cap U}(0)\cap U=N_{C-D}^{U}(0),$ (13)
where Equation 11 holds by definition of nested normals, Equation 12 follows from Lemma 6.6 applied to the sets $C\cap(U+S)$ and $D\cap(U+S)$, and Equation 13 follows from point 3) of $A(U)$. This shows that point 2) of $A(U)$ holds for $U$, and the proof is complete. ∎
We now have the tools to prove Corollary 4.1.
Proof of Corollary 4.1.
By Definition 6.5, $T(C,D)=H_{C-D}(0)$. So $T\in\mathcal{F}_{C-D}(0)$, which by Theorem 6.4 implies that, taking $S=C\cap D$, $T\in\mathcal{F}_{C}(C\cap D)\cap\mathcal{F}_{D}(C\cap D)$. By the definition of supporting subspaces (Definition 6.2), $C\cap D-C\cap D\subseteq T$. ∎
7 Proofs
We begin with the proof of Proposition 5.1, after which each subsection is dedicated to one result from Section 4.
Proof of Proposition 5.1.
First we show that for $C\subseteq\mathbb{R}^{n}$ convex, and $x\in\operatorname{relint}(C)$, $F_{C}(x)=C$. Note that
$N_{C}^{\operatorname{\mathrm{span}}(C-C)}(x)=N_{C}^{\operatorname{\mathrm{span}}(C-C)}(C)=(C-C)^{\perp}\cap\operatorname{\mathrm{span}}(C-C)=\{0\},$
where the first equality follows by Lemma 6.1 and the second by Lemma 6.2. Then by Theorem 6.3, $H_{C}(x)=\operatorname{\mathrm{span}}(C-C)$, and by Lemma 6.5, $F_{C}(x)=(\operatorname{\mathrm{span}}(C-C)+x)\cap C=C$.
From Lemma 4.1, for $x\in\operatorname{relint}(\operatorname{dom}(f))\cap\operatorname{relint}(\operatorname{dom}(g))$ and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
$T_{a}\cap\operatorname{dom}(f)=F_{\operatorname{dom}(f)}(\operatorname{dom}(f)\cap\operatorname{dom}(g))\supseteq F_{\operatorname{dom}(f)}(x)=\operatorname{dom}(f).$
From this the results follows. ∎
7.1 Regularity in the joint supporting subspace
We now show the main property of the joint supporting subspace, Theorem 4.2. We first state and prove a lemma which connects the notion of separation of convex set to a statement reminiscent of the qualification condition by Mordukhovich in variational analysis: [19, Theorem 2.19].
Lemma 7.1 (Separation from local normal cones).
For convex sets $C,D\subseteq\mathbb{R}^{n}$ and $S\subseteq C\cap D$ nonempty, the following equivalence holds:
$N_{C}(S)\cap(-N_{D}(S))\neq\{0\}\iff C\text{ is separated from }D.$ (14)
Proof of Lemma 7.1.
We start with $(\implies)$. There exists a $v\in[N_{C}(S)\cap(-N_{D}(S))]\setminus\{0\}$, and from Lemma 6.2 $v\perp S-S$, whence the hyperplane $v^{\perp}+S$ separates $C$ from $D$.
[…]
  1. 1.
    $\langle v,s\rangle=\beta$ for all $s\in S$;
  2. 2.
    $\langle v,c\rangle\geq\beta$ for all $c\in C$;
  3. 3.
    $\langle v,d\rangle\leq\beta$ for all $d\in D$.
Thus, for all $y\in C$ and $s\in S$, $\langle v,y-s\rangle\geq 0$, so $v\in N_{C}(S)$. Similarly, $v\in N_{D}(S)$, whence $v\in N_{C}(S)\cap(-N_{D}(S))\neq\{0\}$. ∎
Proof of Theorem 4.2.
Recall that from Definition 6.5, $T=H_{C-D}(0)$. Then by Theorem 6.3, $N_{C-D}^{T}(C\cap D)=\{0\}$. Applying Theorem 6.4, we find that $N^{T}_{C}(C\cap D)\cap(-N^{T}_{D}(C\cap D))=\{0\}$.
By applying Lemma 7.1 in the space $T_{a}$, the result follows. Note that Lemma 7.1 is stated in $\mathbb{R}^{n}$, yet we can apply it in $T_{a}$ because $T$ is a finite-dimensional subspace, thereby isomorphic to $\mathbb{R}^{n}$, and Lemma 7.1 is invariant when translating $T$ to $T_{a}$. ∎
7.2 Analytical characterizations
In this section we prove Lemma 4.1 and Theorem 4.1.
Lemma 7.2 (Nested normals of constrained set).
Let $C\subseteq\mathbb{R}^{n}$ be a convex set and $S\subseteq C$ a nonempty set. Let $U,V\subseteq\mathbb{R}^{n}$ be subspaces such that $S-S\subseteq U\subseteq V$ and $C\cap(V+S)\subseteq U+S$. Then
$N_{C}^{U}(S)+(U^{\perp}\cap V)=N_{C\cap(U+S)}^{V}(S).$
Proof of Lemma 7.2.
We begin with the containment ($\supseteq$). Let $x\in N_{C}^{V}(S).$ Then $\forall s\in S,y\in C\cap(V+S)$, note that $y-s\in U$, and so
$\langle\operatorname{\Pi}_{U}x,y-s\rangle=\langle x,y-s\rangle\leq 0.$
This shows that $\operatorname{\Pi}_{U}x\in N_{C}^{U}(S)$. Now $x=\operatorname{\Pi}_{U}x+\operatorname{\Pi}_{U^{\perp}}x$, and $x,\operatorname{\Pi}_{U}x\in V\implies\operatorname{\Pi}_{U^{\perp}}x\in V$, therefore $x\in N_{C}^{U}(S)+(U^{\perp}\cap V)$.
For the inclusion ($\subseteq$), we consider any $q\in N^{U}_{C}(S)$ and $z\in U^{\perp}\cap V$. Let $y\in C\cap(U+S),s\in S$, and note that $\langle z,y-s\rangle=0$ because $y-s\in U$. Then
$\langle q+z,y-s\rangle=\langle q,y-s\rangle\leq 0.$
Therefore $q+z\in N_{C\cap(U+S)}^{V}(S)$, which shows the inclusion and completes the proof. ∎
Lemma 7.3 (Conical projections yield supporting subspaces).
Let $C\subseteq\mathbb{R}^{n}$ be convex, $S\subseteq C$ non-empty, and $U\in\mathcal{F}(C;S)$. Then for any $K\subseteq N^{U}_{C}(S)$,
$U\cap K^{\perp}\in\mathcal{F}(C;S).$
Proof of Lemma 7.3.
Since $K$ is finite dimensional, there is a finite number of vectors $v_{1},\ldots,v_{m}\in K$ such that $\operatorname{\mathrm{span}}(K)=\operatorname{\mathrm{span}}(v_{1},\ldots,v_{m})$. Then since $v_{1},\ldots,v_{m}\in N_{C}^{U}(S)$, by repeated application of Theorem 6.1, $U\cap v_{1}^{\perp}\ldots\cap v_{m}^{\perp}$ is a supporting subspace, which is the same set as $U\cap K^{\perp}$. ∎
Lemma 7.4 (Generated supporting subspace of one of the two convex sets).
For convex sets $C,D\subseteq\mathbb{R}^{n}$, $T=T(C,D)$, $T_{a}=T+C\cap D$, we have
$H_{C}(C\cap D)=T\cap N_{C}^{T}(C\cap T_{a})^{\perp}.$
Proof of Lemma 7.4.
Let $U=T\cap N_{C}^{T}(C\cap T_{a})^{\perp}$. we first show that $U\in\mathcal{F}(C\cap D;C)$. Note that $T\in\mathcal{F}(C\cap D;C)$ by Theorem 6.4. Then we can use Lemma 7.3 because $N_{C}^{T}(C\cap T_{a})\subseteq N_{C}^{T}(C\cap D)$, which shows that $U\in\mathcal{F}(C\cap D;C)$ as well.
In the remainder of the proof we show that the assumption of Theorem 6.3 holds for $U$, namely that $N_{C}^{U}(C\cap D)=\{0\}$.
We first show that $N_{C}^{U}(C\cap T_{a})=\{0\}$. By Lemma 6.2, $N_{C}^{T}(C\cap T_{a})=(C\cap T_{a}-C\cap T_{a})^{\perp}\cap T$. Also, note that by the identity in Lemma 7.2, we find that $N_{C}^{U}(C\cap T_{a})=N_{C}^{T}(C\cap T_{a})\cap U$. Combining these two facts,
$N_{C}^{U}(C\cap T_{a})=N_{C}^{T}(C\cap T_{a})\cap U=(C\cap T_{a}-C\cap T_{a})^{\perp}\cap U.$
But note that $C\cap T_{a}-C\cap T_{a}\subseteq U$, therefore it fo $N_{C}^{U}(C\cap T_{a})=\{0\}$. That $C\cap T_{a}-C\cap T_{a}\subseteq U$ holds because $C\cap T_{a}-C\cap T_{a}\subseteq T$ and $(C\cap T_{a}-C\cap T_{a})\perp N_{C}^{T}(C\cap T_{a})$ by Lemma 6.2.
It only remains to show that $N_{C}^{U}(C\cap D)=N_{C}^{U}(C\cap T_{a})$. By Corollary 4.5, we may pick some $x\in\operatorname{relint}(C\cap T_{a})\cap\operatorname{relint}(D\cap T_{a})$. Then by Lemma 6.1,
$N_{C}^{U}(C\cap T_{a})=N_{C}^{U}(x)\supseteq N_{C}^{U}(C\cap D).$
Then $N_{C}^{U}(C\cap D)=\{0\}$, and so the result follows from Theorem 6.3. ∎
We are now in a position to prove the analytical characterizations in Section 4.
Proof of Lemma 4.1.
From Lemma 7.4, $H_{C}(C\cap D)=T\cap N_{C}^{T}(C\cap T_{a})^{\perp}$. Since $(C\cap T_{a}-C\cap T_{a})\perp N_{C}^{T}(C\cap T_{a})$ by Lemma 6.2, $(C\cap T_{a}-C\cap T_{a})\subseteq H_{C}(C\cap D)$, i.e., $H_{C}(C\cap D)+C\cap T_{a}\supseteq C\cap T_{a}$. But $H_{C}(C\cap D)+C\cap T_{a}=H_{C}(C\cap D)+C\cap D$, therefore
$H_{C}(C\cap D)+C\cap D\supseteq C\cap T_{a}.$ (15)
Additionally, since we know that $T\in\mathcal{F}(C\cap D;C)$ from Theorem 6.4, by definition of the generated supporting subspace we know that $H_{C}(C\cap D)\subseteq T$. Therefore,
$H_{C}(C\cap D)+C\cap D\subseteq T_{a}.$ (16)
From Equation 15 and Equation 16 it follows that $(H_{C}(C\cap D)+C\cap D)\cap C=C\cap T_{a}$. Then the statement follows from Lemma 6.5. ∎
Proof of Theorem 4.1.
From Lemma 4.1, we have that $T\in\mathcal{F}(C;C\cap D)$ is a supporting subspace of $C$ that contains $C\cap D$, and so by Definition 6.4, $H_{C}(C\cap D)\subseteq T$. Therefore, for $^{\prime}:=H_{C}(C\cap D)+H_{D}(C\cap D), TT$$^{\prime}\subseteq T$.
For the other containment, we argue by contradiction: let $T^{\prime}\subsetneq T$. Then there is a unit vector $v\in T$ with $v\perp T^{\prime}$. Then $v\perp H_{C}(C\cap D)$, which by Lemma 7.4, implies that $v\in N_{C}^{T}(C\cap T)$. By Lemma 6.2, $v\in N_{C}^{T}(C\cap T)$ as well. And since $N_{C}^{T}(C\cap T)\subseteq N_{C}^{T}(C\cap D)$, we find that $v,-v\in N_{C}^{T}(C\cap D)$. By a similar argument, $v,-v\in N_{D}^{T}(C\cap D)$. But then $v\in N_{C}^{T}(C\cap D)\cap(-N_{D}^{T}(C\cap D))$, which by Theorem 6.4 means that $v\in N_{C-D}^{T}(0)$. But by Theorem 6.3, $T$ cannot contain such a nested normal because by Definition 6.5, $T=H_{C-D}(0)$. Therefore, by contradiction $T=T^{\prime}$. ∎
Next, we present the proofs for the corollaries of Theorem 4.1, starting with a simple lemma.
Lemma 7.5 (Formulation of the tangent space).
For a convex set $C\subseteq\mathbb{R}^{n}$ and $S\subseteq C$ non-empty,
$\operatorname{\mathrm{span}}(C-C)=\operatorname{\mathrm{span}}(C-S).$
Proof of Lemma 7.5.
That $\operatorname{\mathrm{span}}(C-C)\supseteq\operatorname{\mathrm{span}}(C-S)$ is trivial, so we only show $\operatorname{\mathrm{span}}(C-C)\subseteq\operatorname{\mathrm{span}}(C-S)$. It suffices to show that $C-C\subseteq\operatorname{\mathrm{span}}(C-S)$; let $u,v\in C$, then for any fixed $x\in S$,
$u-v=(u-x)-(v-x)\in\operatorname{\mathrm{span}}(C-S).$
Proof of Corollary 4.3.
In this proof, we denote $F_{C}:=F_{C}(C\cap D)$ and $F_{D}:=F_{D}(C\cap D)$.
$\operatorname{\mathrm{span}}(F_{C}-F_{D})$ $=\operatorname{\mathrm{span}}(F_{C}-F_{D})+(C\cap D-C\cap D)$ (17)
$=\operatorname{\mathrm{span}}(F_{C}-F_{D}+(C\cap D-C\cap D))$ (18)
$=\operatorname{\mathrm{span}}((F_{C}-C\cap D)-(F_{D}-C\cap D))$ (19)
$=\operatorname{\mathrm{span}}(F_{C}-C\cap D)+\operatorname{\mathrm{span}}(F_{D}-C\cap D)$ (20)
$=\operatorname{\mathrm{span}}(F_{C}-F_{C})+\operatorname{\mathrm{span}}(F_{D}-F_{D}),$ (21)
$=H_{C}(C\cap D)+H_{D}(C\cap D).$ (22)
where Equation 17 holds because
$(C\cap D-C\cap D)\subseteq F_{C}-F_{D}\subseteq\operatorname{\mathrm{span}}(F_{C}-F_{D}),$
and Equation 21 follows from Lemma 7.5. The result then follows from Theorem 4.1. ∎
Proof of Corollary 4.2.
Throughout this proof, we omit the argument $C\cap D$, i.e., $H_{C}:=H_{C}(C\cap D)$, and similarly for $H_{D},F_{C},F_{D}$. We also define $\bar{F}_{C}=F_{C}-C\cap D$ and $\bar{F}_{D}=F_{D}-C\cap D$. By Lemma 7.5,
$H_{C}=\operatorname{\mathrm{span}}(F_{C}-F_{C})=\operatorname{\mathrm{span}}(\bar{F}_{C}).$
With this, starting with $T_{a}$ as in Theorem 4.1,
$T_{a}(C,D)=H_{C}+H_{D}+C\cap D$
$=\operatorname{\mathrm{span}}(\bar{F}_{C})+\operatorname{\mathrm{span}}(\bar{F}_{D})+C\cap D$
$=\operatorname{\mathrm{span}}(\operatorname{\mathrm{span}}(\bar{F}_{C})\cup\operatorname{\mathrm{span}}(\bar{F}_{D}))+C\cap D$
$=\operatorname{\mathrm{span}}(\bar{F}_{C}\cup\bar{F}_{D})+C\cap D=\operatorname{aff}(\bar{F}_{C}\cup\bar{F}_{D})+C\cap D$
$=\operatorname{aff}(\bar{F}_{C}\cup\bar{F}_{D}+C\cap D)=\operatorname{aff}(F_{C}\cup F_{D}).$
7.3 Second characterization
We can now state the second characterization in full.
Theorem 7.1 (Iterative bilateral facial reduction).
Let $C,D\subseteq\mathbb{R}^{n}$ be convex sets and $S\subseteq C\cap D$ non-empty. Beginning with $T_{0}=\mathbb{R}^{n}$, let
$T_{i+1}=T_{i}\cap(N^{T_{i}}_{C}(S)\cap[-N^{T_{i}}_{D}(S)])^{\perp}.$
Let $\ell$ be the first index for which $T_{\ell+1}=T_{\ell}$. Then:
  1. 1.
    $T_{\ell}=T(C,D)$.
  2. 2.
    $\ell\leq n$.
Proof of Theorem 7.1.
First, $(N^{T_{i}}_{C}(S)\cap[-N^{T_{i}}_{D}(S)])=N^{T_{i}}_{C-D}(S)$ by Theorem 6.4. Applying Lemma 7.3 at each step, we find that $\forall i\in\{0,\ldots,\ell\},\,T_{i}\in\mathcal{F}(C-D;S)$. Finally, note that whenever $N^{T_{i}}_{C}(S)\cap[-N^{T_{i}}_{D}(S)]\neq\{0\}$, $\dim(T_{i+1})\leq\dim(T_{i})-1$. Therefore, the dimension may be reduced at most $n$ times, which implies that $\ell\leq n$. On the other hand, the fact that $T_{\ell}=T_{\ell+1}$ implies that $N^{T_{\ell}}_{C-D}(S)=\{0\}$. Then from Theorem 6.3 it follows that $T_{\ell}=H_{C-D}(S)$, which is $T(C,D)$ from Definition 6.5. ∎
Proof of Corollary 4.4.
We consider the formula
$T_{i+1}=T_{i}\cap(N^{T_{i}}_{C}(x)\cap[-N^{T_{i}}_{D}(x)])^{\perp}.$
from Theorem 7.1. We show that the r.h.s. is the same set as
$T_{i}\cap(N_{C\cap(T_{i}+x)}(x)\cap[-N_{D\cap(T_{i}+x)}(x)])^{\perp}.$
We use a substitution as in Lemma 7.2 with $V=\mathbb{R}^{n}$.
$T_{i}\cap(N_{C\cap(T_{i}+x)}(x)\cap[-N_{D\cap(T_{i}+x)}(x)])^{\perp}$
$=$ $T_{i}\cap((N^{T_{i}}_{C}(x)+T_{i}^{\perp})\cap[-N^{T_{i}}_{D}(x)+T_{i}^{\perp}])^{\perp}$
$=$ $T_{i}\cap(N^{T_{i}}_{C}(x)\cap[-N^{T_{i}}_{D}(x)]+T_{i}^{\perp})^{\perp}$
$=$ $T_{i}\cap(N^{T_{i}}_{C}(x)\cap[-N^{T_{i}}_{D}(x)])^{\perp}.$
The last equality holds because $N^{T_{i}}_{C}(x)\cap[-N^{T_{i}}_{D}(x)]\subseteq T_{i}$, so we may use the identity that $T_{i}\cap(S+T_{i}^{\perp})^{\perp}=T_{i}\cap S^{\perp}$ for any set $S\subseteq T_{i}$. ∎
7.4 Conclusion
In this work we define and characterize the joint supporting subspace $T_{a}$, which exists between any two convex sets. By defining this affine subspace from the domains of two convex functions, and then reducing each function to it, qualification conditions are guaranteed to hold. We obtain a number of characterizations for $T_{a}$ by bridging between the facial structure of the difference of two convex sets CDC-D and the facial structure of the individual convex sets. By doing this, we show that “pathology on the boundary” of convex functions can be tamed by considering the facial structure of their domains.
Future work avenues include specifying and implementing efficient reduction algorithms for convex functions, and pairing them with optimization solvers to obtain robust algorithms with strong theoretical guarantees.
7.5 Acknowledgements
The author would like to thank Sharvaj Kubal, Naomi Graham, prof. Yaniv Plan, and prof. Ozgur Yilmaz for many fruitful discussions. A special thanks to prof. Michael Friedlander for pointing out that a generalization of the Fenchel-Rockafellar dual and optimality conditions would be a promising research direction. This paper was produced with the Latex Exporter plugin in Obsidian.
Declarations
The author is supported by the University of British Columbia.